### v3

# Uncertainty Quantification on Graph Learning: A Survey

Abstract
Graphical models have demonstrated their exceptional capabilities across numerous applications, such as social networks, citation networks, and online recommendation systems. However, their performance, confidence, and trustworthiness are often limited by the inherent randomness in data and the challenges of accurately modeling real-world complexities. There has been increased interest in developing uncertainty quantification (UQ) techniques tailored to graphical models. In this survey, we comprehensively examine existing works on UQ for graphical models, focusing on key aspects such as the sources, representation, handling, and evaluation of uncertainty. This survey distinguishes itself from most existing UQ surveys by specifically concentrating on UQ in graphical models, including probabilistic graphical models (PGMs) and graph neural networks (GNNs). After reviewing sources of uncertainty, we organize the work using two high-level dimensions: uncertainty representation and uncertainty handling. By offering a comprehensive overview of the current landscape, including both established methodologies and emerging trends, we aim to bridge gaps in understanding key challenges and opportunities in UQ for graphical models, hoping to inspire researchers working on graphical models or uncertainty quantification to make further advancements at the cross of the two fields.
Index Terms:
Article submission, IEEE, IEEEtran, journal, , paper, template, typesetting.
I Introduction
Graphical models have emerged as fundamental tools in machine learning due to their capability to capture and manage complex relational dynamics among variables. These models have found extensive applications, such as spam detection in online review systems [1, 2], relationship extraction in social networks [3, 4, 5, 6], user interest exploration in recommendation systems [7, 8, 9, 10], and drug discovery in protein-protein interaction networks [11, 12, 13]. Unlike traditional models, where variables are assumed to be independently and identically distributed (i.i.d.), graphical models could reveal intricate relationships and conditional dependencies. By considering both individual variables and their interconnections, graphical models improve predictive performance in complex data environments and tasks.
However, uncertainties are ubiquitous and can arise from various sources, including inherent data randomness, model training errors, and unexpected test data distributions [14]. These uncertainties hinder the deployment of graphical models in high-stakes domains where precise risk assessment and decision-making under uncertainty are vital, including healthcare [15, 16] and autonomous driving [17, 18]. Uncertainty quantification (UQ) provides calibrated confidence measures alongside predictions, thereby improving the reliability of the models in practice [19, 20, 21]. Thus, many studies focus on developing UQ techniques for graphical models, enabling reliable applications in critical fields such as climate risk forecasting [22] and intelligent control systems [23].
This survey explores various types of uncertainties in two primary types of graphical models: probabilistic graphical models (PGMs) and graph neural networks (GNNs). PGMs, including Bayesian Networks and Markov Random Fields, utilize more rigorous probabilistic formulations to effectively represent uncertainty and are particularly effective in scenarios requiring joint probability modeling and causal relationship analysis. GNNs extend the successes of deep learning to graph-structured data but face uncertainty challenges such as noisy links, missing nodes, or mislabeled data. Recent developments have incorporated uncertainty quantification into the architectures and/or training algorithms of GNNs, improving their robustness and interoperability.
Figure 1: We first show an example from a recommendation system, where the user of interest (in gray) links to other users and contents with various relationships. A model’s predictions for the user’s preference under uncertainty are shown below, with or without uncertainty handling methods. Then, as detailed in Sec. III, the source of uncertainty can be decomposed into aleatoric and epistemic uncertainties. Last, we demonstrate the state-of-the-art methods proposed for uncertainty representation and handling, respectively. Specifically, uncertainty representation methods (refer to Sec. IV) include graph variational autoencoder, Bayesian graph neural network, and OOD epistemic uncertainty representation methods. Besides, as discussed in Sec. V, uncertainty handling methods such as OOD handling techniques, conformalization, and calibration, are shown in the figure.
Existing surveys on UQ provide comprehensive coverage of UQ methods, and offer valuable insights into UQ across various application domains. These surveys can be categorized based on specific application domains (e.g., UQ for graph learning) and the extent of their coverage, ranging from broad overviews of UQ methods [24, 25, 26, 27] to focused studies on particular areas such as scientific machine learning [28], sources of uncertainty [29, 30], optimization under uncertainty [31], and noise in machine learning [32]. Other surveys focus on specific paradigms, like integrating differential equations (DEs) with GNNs, which includes some discussion of uncertainty within the DE context [33]. However, none of these focus on uncertainty quantification on graph learning.
The survey most closely related to ours is [14], which provides a broad overview of UQ in graph learning. However, Wang et al. [14] focus primarily on downstream tasks and applications related to GNN uncertainty, rather than on the theoretical foundations of UQ methods themselves. In contrast, our work delves into the mathematical underpinnings of each UQ approach, presenting the key formulas and core concepts that offer a principled perspective on uncertainty in graphs. Moreover, whereas Wang et al. categorize methods by “single models versus ensemble models” and “deterministic models versus models with random parameters,” these labels obscure the conceptual differences between approaches. We instead organize UQ methods into “Uncertainty representation” (how uncertainty is modeled) and “Uncertainty handling” (how uncertainty is utilized or mitigated), which more clearly reflects the roles of different techniques. Finally, our survey incorporates advanced UQ for graphs methodologies, such as frameworks based on stochastic differential equations, which are entirely absent from [14].
Fig. 1 shows an example of uncertainty on a recommendation system, and highlights three key components of the survey: source, representation, and handling of uncertainties. For the organization, we first provide the preliminary knowledge concerning PGMs and GNNs in Sec. II, and introduce the sources of uncertainty in Sec. III. We then discuss the existing studies for uncertainty representation and uncertainty handling in Sec. IV and Sec. V, respectively. Furthermore, we demonstrate widely used metrics for uncertainty evaluation in Sec. VI. Finally, future directions and the conclusion of the survey are discussed in Sec. VII and Sec. VIII, respectively.
II Preliminaries
TABLE I: Notation Definitions
$G=(\mathcal{V},\mathcal{E})$ A graph with node set $\mathcal{V}$ and edge set $\mathcal{E}$.
$V_{i}\in\mathcal{V}$ A node with features in the graph.
$X_{i}$ Node feature vector of node $V_{i}$.
$Y_{i}\in\mathcal{Y}$ True label of node $V_{i}$.
$\hat{Y}_{i}\in\mathcal{Y}$ Predicted label of node $V_{i}$.
$K$ The number of classes in a classification task.
$\mathcal{N}(V_{i})$ Direct neighbors of node $V_{i}$.
$f(\boldsymbol{\theta})\in\mathcal{F}$ A model $f$ with parameters $\boldsymbol{\theta}$ from a function space $\mathcal{F}$.
$P_{\ast}$, $P_{0}$ Real and observed data distribution.
$x_{i},y_{i}$ A specific value of $X_{i}$ and $Y_{i}$, respectively.
II-A Basic Notations for Graphs
We summarize the key notations in Table I. Consider a graph $G=(\mathcal{V},\mathcal{E})$ consisting of a set of $n$ nodes $\mathcal{V}=\{V_{1},\dots,V_{n}\}$, each representing a random variable, and each variable $V_{i}\in\mathcal{V}$ is associated with a feature vector $X_{i}$. The “feature” could be a prior distribution indicating the likelihood of belonging to class $k$ (typically in MRF), or a vector with practical meanings, such as bag-of-words (typically in GNN). Each edge in the set of edges $\mathcal{E}$ is a relationship between two variables. The set $\mathcal{N}(V_{i})$ collects direct neighbors of $V_{i}$, e.g., $\mathcal{N}(V_{i})=\{V_{j}:(V_{i},V_{j})\in\mathcal{E}\}$.
II-B Probabilistic Graphical Model
A probabilistic graphical model (PGM) [34, 35] is widely used to represent the joint probability distribution of random variables, whose relations can be depicted in a graph. There are two distinct categories of PGMs, depending on whether the graph’s edges are directed or undirected. Specifically, a Bayesian Network (BN) uses a directed graph, where edges represent causality. Conversely, a Markov Random Field (MRF) uses an undirected graph, where edges signify the correlation between nodes. Both BN and MRF factorizes the joint distribution $P(\mathcal{V})$ as follows [35]:
$P(\mathcal{V})=\frac{1}{Z}\prod_{s\in\mathcal{S}}\zeta_{s}(X_{s}),$ (1)
where each factor $\zeta_{s}(X_{s})$ is a nonnegative function defined on a subset $X_{s}\subseteq\{X_{1},X_{2},\dots,X_{n}\}$, and $Z$ is for normalization. A BN defines $\zeta_{s}$ as the node’s probability distribution conditioned on their parents and naturally $Z=1$. Formally,
$P(\mathcal{V})=\prod_{i}^{n}P(X_{i}|Pa(X_{i})),$ (2)
where $Pa(X_{i})$ specifies parent nodes of $X_{i}$. An MRF factorizes the joint distribution $P(\mathcal{V})$ as
$P(\mathcal{V})=\frac{1}{Z}\prod_{V_{i}\in\mathcal{V}}\phi(X_{i}=x_{i})\prod_{V_{j}\in\mathcal{N}(V_{i})}\psi(X_{i}=x_{i},X_{j}=x_{j}),$ (3)
where $Z$ normalizes the product to a probability distribution. $\phi(X_{i}=x_{i})$ denotes the prior probability of $X_{i}$ taking value $x_{i}$, independent of other variables. The compatibility $\psi(X_{i}=x_{i},X_{j}=x_{j})$ represents the probability of the pair $(x_{i},x_{j})$ occurring jointly, capturing the dependencies between the two variables. The marginal distribution $b(X_{i})$ for $V_{i}\in\mathcal{V}$ is
$b(X_{i}=x_{i})=\sum_{\mathcal{V}\setminus\{X_{i}\}}P(\mathcal{V}\setminus\{X_{i}\},X_{i}=x_{i}).$ (4)
II-C Graph Neural Networks
Unlike traditional neural networks that operate on grid-like data (such as tabular vectors, images, or sequences), graph neural network (GNN) could learn from complex relationships in graphs. Formally, for the $l$-th layer of an $L$-layers GNN, the representation $\mathbf{h}_{i}^{(l)}$ of node $V_{i}$ is
$\mathbf{a}^{(l)}_{i}$ $=$ $\textnormal{AGG}\left(\left\{\mathbf{h}^{(l-1)}_{j}|V_{j}\in\mathcal{N}(V_{i})\right\};\theta^{(l)}\right),$ (5)
$\mathbf{h}^{(l)}_{i}$ $=$ $\textnormal{UPDATE}\left(\mathbf{a}^{(l)}_{i},\mathbf{h}^{(l-1)}_{i};\theta^{(l)}\right).$ (6)
The AGG function aggregates the messages sent from all neighbors $V_{j}\in\mathcal{N}(V_{i})$ to $V_{i}$, and it could operate as an element-wise average, or maximum over the message set, or employ a parametric layer for aggregation [36]. The UPDATE function combines representations of node $X_{i}$ and its neighbors. $\mathbf{h}^{(0)}_{i}$ is the input node feature for node $X_{i}$. The representations from the last layer $\mathbf{h}_{i}^{(L)}$ can be softmaxed to a probability distribution for node-classification tasks [6], used in a scoring function for link prediction tasks [37], or mapped to a graph representation by a READOUT function for graph-level tasks. The parameters of GNN, $\boldsymbol{\theta}=\{\theta^{(l)}$: $l=1,\dots,L$, are trained in an end-to-end fashion.
III Sources of Uncertainty
The total uncertainty in machine learning can be decomposed into aleatoric uncertainty (AU) and epistemic uncertainty (EU) based on their sources [38]. [^1] Recent studies have challenged the traditional binary classification of uncertainty into aleatoric and epistemic types, suggesting that this dichotomy may be overly simplistic [39, 40]. Aleatoric uncertainty, which is irreducible with increasing amount of data, represents the randomness inherent to the data generation process [41]. On the other hand, epistemic uncertainty can be reduced by obtaining more knowledge or data [42, 30]. Techniques for the disentanglement of aleatoric and epistemic uncertainty are introduced by Valdenegro-Toro et al. [43], and Munikoti et al. [44] formulate mathematical expressions of these two kinds of uncertainty in the context of GNNs.
The identification of aleatoric and epistemic uncertainty must be built upon the context of the selected model and available data [45]. For example, flipping a coin is usually regarded as a purely random event, involving aleatoric uncertainty. However, suppose the entire physics process of flipping a coin is known, including the flipping force, the coin’s mass, volume, initial position, etc.. In that case, the process can be simulated, and whether the result would be a head or a tail is predictable. With such a physical simulation, randomness can be regarded as epistemic uncertainty and can be reduced or even eliminated.
To reduce the ambiguity mentioned above, we mathematically define the two sources of uncertainty and how they impact prediction reliability. Let $X\in\mathcal{X}\subseteq\mathbb{R}^{d}$ and $Y\in\mathcal{Y}\subseteq\mathbb{R}$ denote the input and output variables with realizations $x$ and $y$, respectively. The underlying data distribution is $P_{XY}$ and the data generation mechanism is $f_{0}$ such that $y=f_{0}(x)$.
III-A Aleatoric Uncertainty
Formally, when the data-generating mechanism $f_{0}$ is conditioned on $X=x$, aleatoric uncertainty (AU) can be represented by the conditional distribution $P_{Y|X}(\cdot\mid x)$. The dispersion of this distribution quantifies the AU and is often measured by the conditional Variance $\mathrm{\text{Var}}_{Y|X}[Y\mid x]\in\mathbb{R}$. . Gruber et al. [29] identify four distinct sources of aleatoric uncertainty, categorized as follows.
Omitted Features. Omitted features are variables excluded during model training and prediction. They contribute to AU, since collecting more data without these missing features does not reduce the uncertainty they introduce. The omission is investigated in the context of causal models [46], confounder analysis [47], and model sensitivity analysis [48]. The omission may result from variables being unobservable or overlooked during data collection. Let $Z\in\mathcal{Z}$ denote an unobserved variable that influences the value of $Y$, the expectation of $Y|x$ is a weighted integration of $\mathbb{E}_{Y|X,Z}$ on $Z$:
[…]
Gruber et al. [29] also point out that $\text{Var}_{Y|X,Z}[Y|x,z]\leq\text{Var}_{Y|X}[Y|x]$ with at least one $z\in\mathcal{Z}$, indicating that including the non-negligible variables will reduce AU.
Feature Errors. Feature errors represent inaccurate feature values in collected data, which are likely to occur with improper measurements, such as low-resolution cameras, low-quality audio recorders, and uncalibrated gauges [32]. These errors can cause biases and increased Variance in feature collection. Let $Z$ represent feature variables by faultless measurements and $X$ denote those by inaccurate measurements, assuming that $Y$ is independent of $X|Z$. The AU caused by feature errors is illustrated as
$\text{Var}_{Y|X}[Y|x]=$ $\mathbb{E}_{Z|X}\left[\text{Var}_{Y|Z}[Y|Z]|x\right]$ (8)
$\;+\text{Var}_{Z|X}\left[\mathbb{E}_{Y|Z}[Y|Z]|x\right],$
where the first term of the right-hand side approximates $\text{Var}_{Y|Z}(Y|Z)$ if $Y\perp\!\!\!\perp X|Z$, and the second term is an increment in AU caused by feature errors [29].
[…]
This decomposition shows that the observed label distribution $P_{Y|X}$ can be viewed as a mixture over the true label distribution $P_{Z|X}$, modulated by the conditional noise model $P_{Y|X,Z}$, which captures the randomness introduced by the labeling process. In this case, uncertainty in $Y$ arises not only from inherent ambiguity in the input–output relationship, but also from label noise. As such, this contributes to aleatoric uncertainty, since it cannot be reduced merely by increasing the number of training samples without improving label quality.
Missing Data. Missing data indicates some entries in the dataset are incomplete [50]. This differs from omitted features, which refer to entire variables that are excluded from the dataset, rather than individual missing values within an included feature. Denoting $(x_{i},y_{i})$ the $i$-th sample of the dataset, $M_{i}$ is defined as the corresponding missing indicator of $(x_{i},y_{i})$. Specifically, $M_{i}=0$ means the sample is complete, and $M_{i}=1$ otherwise. With a bias factor ${\Pr(M=0|y,x)}/{\Pr(M=0|x)}$ [29], the influence from incompleteness can be expressed by
$P_{Y|X,M}(y|x,M=0)=\frac{\text{Pr}(M=0|y,x)}{\text{Pr}(M=0|x)}P_{Y|X}(y|x).$ (10)
This equation reveals that missing data can introduce selection bias into the observed conditional distribution $P_{Y\mid X,M}$, which deviates from the true distribution $P_{Y\mid X}$. The uncertainty arising from this bias is aleatoric in nature, as it stems from incomplete observations that cannot be resolved simply by collecting more data unless the missingness mechanism is addressed or corrected.
III-B Epistemic Uncertainty
Epistemic uncertainty (EU) is from model selection and the lack of data and knowledge. EU can be categorized into model, approximation, and Out-of-Distribution (OOD) uncertainties [30].
Model Uncertainty. Model uncertainty refers to the difference between the ground truth mapping $f_{0}\in\mathcal{F}_{0}$ and the best possible predictor $f\in\mathcal{F}$. It is impossible to explore the entire function space $\mathcal{F}_{0}$, resulting in a gap between $f$ and $f_{0}$. Specifically, taking $l$ as the loss function during training, $f_{0}$ and $f$ can be explicitly defined as
[…]
Thereby, we can represent model uncertainty by $D(P_{XY},(\mathrm{id},f)_{\#}P_{X})$, where $D(\cdot,\cdot)$ is a user-specified discrepancy measure and $(\mathrm{id},f)_{\#}P_{X}$ is the pushforward measure of $P_{X}$ through the mapping $x\mapsto(x,f(x))$, i.e., the joint distribution of $(x,f(x))$ when $x\sim P_{X}$..
Approximation Uncertainty. Approximation uncertainty is defined as the uncertainty due to the limited number of $N$ training samples. In other words, $P_{XY}$, as well as $f$ calculated in Eq. (12) is not accessible in practice. Therefore, the induced predictor trained with finite samples is defined as
$\hat{f}=\arg\min_{f\in\mathcal{F}}\frac{1}{N}\sum_{i=1}^{N}l(y_{i},f(x_{i})),$ (13)
which allows us to quantify approximation uncertainty as $D(P_{XY},(\mathrm{id},\hat{f})_{\#}P_{X})$.
OOD Uncertainty. Unlike the previous two types, OOD (Out-of-Distribution) uncertainty arises at the inference stage when the deployed model encounters test data drawn from a distribution different from the training distribution [51, 52]. Formally, let $P_{XY}$ and $Q_{XY}$ denote the training and test distributions, respectively. OOD uncertainty occurs when $P_{XY}\neq Q_{XY}$.
This distributional shift can arise in real-world scenarios, such as changes in data collection environments (e.g., sensor drift in medical devices), evolving user behavior over time (e.g., preferences in recommender systems), or deployment in a new domain (e.g., applying a model trained on urban traffic data to rural areas). Since the model ${f}$ is trained to minimize the expected loss under $P_{XY}$ in Eq. (12), its performance and confidence estimates may be unreliable when evaluated under $Q_{XY}$. Therefore, OOD uncertainty can be formally quantified by $D(Q_{XY},(\mathrm{id},f)_{\#}Q_{X})$.
Model and approximation uncertainties are reducible by expanding the search space in the function space $\mathcal{F}$ and increasing the sample size $N$ [53, 54, 55]. In contrast, OOD uncertainty arises from distributional shifts between training and test data and cannot be mitigated by more data or a larger model alone. Instead, it requires robust training strategies, which will be introduced in Sec. V, to improve generalization under shifted conditions.
IV Methods for Uncertainty Representation
TABLE II: Summary of Uncertainty Representation studies on graphs [^2] Reg., Cla., and Clu. are for Regression, Classification, and Clustering, respectively. NC and GC are for Node and Graph Classification, respectively. NRL and GRL are for Node and Graph Representation Learning, respectively. LP is for Link Prediction. STF is for Spatio-Temporal Forecasting. GG is for Graph Generation..
Methods Task Input Space Model Space
Node Feat. Graph Struc. GNN Msg. Model Params Latent Rep.
Bayesian Representation Learning
AEVB [56] DRL
BNNs [57] Reg. & Cla.
VGAE [58] NRL
GraphVAE [59] GRL
BGCNN [60] NC
BUP [61] NC
GKDE [62] NC
BGNN [44] NC
VGCN [63] NC
Gaussian Process
GGP [64] NC
UaGGP [65] NC
LNK [66] GRL
OOD Epistemic Uncertainty Representation
ADNCE [67] GC
WDRO [68] Clu.
DRGL [69] NC
PR-BCD [70] NC
Xu et al. [71] NC
DWNS [72] NC
RGIB [73] GC
GCC [74] NC & GC
Hu et al. [75] GC
Stochastic Differential Equations
LGNSDE [76] NC
GNSD [77] NC
DVGNN [78] STF
GraphGDP [79] GG
AGGDN [80] STF
TABLE III: Summary of Uncertainty Handling studies on graphs11footnotemark: 1.
Methods Task Post-hoc Non-parametric Input-Adaptive Topology-aware Note
Direct Inference
SocNL [81] NC Label Propagation
GPN [82] NC Evidence Propagation
NETCONF [83] NC Evidence Propagation
Out Of Distribution Techniques
DR-GNN [84] NC Distributionally Robust Optimization
Sadeghi, et al. [85] Reg. Distributionally Robust Optimization
WT-AWP [86] NC Adversarial Training
CAP [87] NC Adversarial Training
GraphDefense [88] NC Adversarial Training
BVAT [89] NC Adversarial Training
GraphAT [90] NC Adversarial Training
GraphCL [91] GC Self-Supervised Learning
SS-GCN [92] NC Self-Supervised Learning
GRACE [93] NC Self-Supervised Learning
Calibration
Histogram Binning [94] Cla. Binning-based
Isotonic Regression [95] Cla. Binning-based
Platt Scaling [96] Cla. Logit Modification
Label Smoothing [97] Cla. Label Modification
Focal Loss [98] Cla. Cross-entropy-based
CaGCN [99] NC Logit Modification
GATS [100] NC Logit Modification
RBS [101] NC Logit Modification
GCL [102] NC Cross-entropy-based
IN-N-OUT [37] LP Logit Modification
E-$\Delta$UQ [103] LP Anchoring-based
G-$\Delta$UQ [104] GC Anchoring-based
Conformal Prediction
Transductive CP [105] Reg. & Cla. Rely on Exchangeability
Inductive CP [106] Reg. & Cla. Rely on Exchangeability
Importance-weighted CP [107] Reg. & Cla. Beyond Exchangeability
CQR [108] Reg. Rely on Exchangeability
APS [109] Cla. Rely on Exchangeability
Robust Validation [110] Reg. & Cla. Beyond Exchangeability
Jurygcn [111] NC Beyond Exchangeability
Transductive DAPS [112] NC Rely on Exchangeability
Inductive NAPS [113] NC Beyond Exchangeability
{forest}
for tree= forked edges, grow’=0, draw=none, rounded corners, node options=align=center,, text width=2.7cm, anchor=center , [Uncertainty Quantification, fill=gray!10 ,rotate=90, parent [Uncertainty Representation Sec. IV, for tree=fill=teal!50, child [Bayesian Representation Learning Sec. IV-A, fill=teal!30, greatgrandchild [Graph Variational Autoencoders: VGAE [58], SIG-VAE [114], DGVAE [115], VGNAE [116], Graphite [117], GraphVAE [59], VGRNN [118], HVGAE [119], fill=teal!10, referenceblock] [Bayesian Graph Neural Networks: Bayesian GCNN [60], BGCN (node copying) [120], BGCN (non-parametric) [121], BUP [61], GKDE [62], VGCN [63], BGNN [44], fill=teal!10, referenceblock]] [Gaussian Process Sec. IV-B, fill=teal!30, greatgrandchild [GGP [64], UaGGP [65], LNK [66], GPGC [122], WGGP [123], GPG [124], DGPG [125], FT/WT-GP [126], GCLGP [127], GGPN [128], fill=teal!10, referenceblock]] [Outlier Representation Sec. IV-C, fill=teal!30, greatgrandchild [Dataset-level uncertainty estimation: ADNCE [67], WDRO [68], DRGL [69] , fill=teal!10, referenceblock] [Sample-level uncertainty representation: CAP [87], Xu et al. [71], DWNS [72], RGIB [73] , fill=teal!10, referenceblock] [Manually designed uncertainty sets: GraphCL [91], Hu et al. [75], GRACE [93] , fill=teal!10, referenceblock] ] [Stochastic Differential Equations Sec. IV-D, fill=teal!30, greatgrandchild [LGNSDE [76], GNSD [77], DVGNN [78], Graphgdp [79], AGGDN [80],fill=teal!10, referenceblock]] ] [Uncertainty Handling Sec. V, for tree=fill=orange!45,child [Bayesian Inference Techniques Sec. V-A, fill=orange!30, greatgrandchild [Exact: SocNL [81], NETCONF [83], GPN [82], fill=orange!10, referenceblock] [Approximate: VGAE [58], SIG-VAE [114], DGVAE [115], VGNAE [116], Graphite [117], GraphVAE [59], VGRNN [118], HVGAE [119],Bayesian GCNN [60], BGCN (node copying) [120], BGCN (non-parametric) [121], GKDE [62], VGCN [63], BGNN [44], GGP [64], UaGGP [65], LNK [66], WGGP [123], DGPG [125], FT/WT-GP [126], GCLGP [127], GGPN [128], LGNSDE [76], DVGNN [78], , fill=orange!10, referenceblock] ] [Out of Distribution Techniques Sec. V-B, fill=orange!30, greatgrandchild [Distributionally Robust Optimization: ADNCE [67], DR-GNN [84], WDRO [68], DRGL [69], Sadeghi, et al. [85] , fill=orange!10, referenceblock] [Adversarial Training: WT-AWP [86], CAP [87] PR-BCD [70], GraphDefense [88], DWNS [72], BVAT [89], GraphAT [90], RGIB [73] , fill=orange!10, referenceblock] [Self-Supervised Learning: GraphCL [91], GCC [74], SS-GCN [92], Hu et al. [75], GRACE [93] , fill=orange!10, referenceblock] ] [Calibration Sec. V-C, fill=orange!30, greatgrandchild [CaGCN [99], GCL [102], MCCNIFTY [129], IN-N-OUT [130], E-$\Delta$UQ [103], G-$\Delta$UQ [104],fill=orange!10, referenceblock]] [Conformal Prediction Sec. V-D, fill=orange!30, greatgrandchild [JuryGCN [111], DAPS [112], NAPS [113],fill=orange!10, referenceblock] ] ] ]
Figure 2: The taxonomy of uncertainty quantification methods in graph learning is organized into two main branches: Uncertainty Representation and Uncertainty Handling. Uncertainty Representation addresses how uncertainty is captured within data and models, including methods such as Bayesian Representation Learning, Gaussian Processes, and Outlier Representation. Uncertainty Handling focuses on methods for evaluating and mitigating uncertainty, encompassing Direct Inference, Out-of-Distribution Techniques, Calibration, and Conformal Prediction. Each leaf node highlights representative works and associated references for each method.
Understanding how uncertainty can be represented is fundamental for designing trustworthy graph learning algorithms. In this section, we delve into several distinct categories of methods for uncertainty representation, including bayesian representation learning, gaussian processes, stochastic differential equations (SDEs) and outlier representation.
IV-A Bayesian Representation Learning
IV-A1 Bayesian Representation Learning Overview
Bayesian methods in representation learning model uncertainty in both model parameters and learned representations. In this survey, we focus on two prominent classes: variational autoencoders (VAEs) [56] and Bayesian neural networks (BNNs)  [131, 132].
VAEs. VAEs introduce uncertainty representation by modeling the data‐generation process with fully probabilistic latent variables rather than a single deterministic code. Concretely, they assume a prior over latent variables $\mathbf{z}\sim p(\mathbf{z})$, commonly a standard normal, and define a likelihood $p_{\boldsymbol{\theta}}(\mathbf{x}\mid\mathbf{z})$, for example a Gaussian distribution whose learned variance captures the aleatoric uncertainty inherent in $\mathbf{x}$. To infer the latent representation for a given input $\mathbf{x}$, the encoder approximates the true posterior with a variational distribution $q_{\boldsymbol{\phi}}(\mathbf{z}\mid\mathbf{x})$, typically a Gaussian whose mean and variance depend on $\mathbf{x}$. Sampling $\mathbf{z}$ from $q_{\boldsymbol{\phi}}(\mathbf{z}\mid\mathbf{x})$ injects stochasticity into the encoding, quantifying uncertainty about which latent variable underlies the observation $\mathbf{x}$.
BNNs. Unlike deterministic neural networks that only provide point estimates, BNNs assign distributions to the parameters $\boldsymbol{\theta}$ to capture model uncertainty. During training, BNNs use Bayesian inference to update the prior distribution $p(\boldsymbol{\theta})$ with the likelihood $p(D|\boldsymbol{\theta})$, producing a posterior distribution that combines prior knowledge with observed data:
$P(\boldsymbol{\theta}|D)=\frac{P(D|\boldsymbol{\theta})P(\boldsymbol{\theta})}{P(D)}.$ (14)
For predictions, BNNs leverage Bayesian model averaging, which involves integrating over the posterior distribution of the parameters to account for uncertainty. For a given test sample $\mathbf{x}^{\ast}$, the prediction incorporates the model uncertainty, yielding the predictive distribution:
[…]
IV-A2 Bayesian Representation Learning for Graphs
The need to represent the complexity of graph-structured data while quantifying uncertainty has driven the development of Bayesian graph representation learning. Unlike deterministic embeddings that provide fixed representations, Bayesian graph representation learning captures the stochastic nature of graphs, accounting for irregular structures, noisy features, and other uncertainties to produce more faithful representations. GNNs lay the groundwork by leveraging local neighborhood information, but their deterministic nature often falls short in capturing the inherent uncertainty of real-world graphs. In response, two prominent approaches have been developed: graph variational autoencoders (GVAEs) [58, 114, 115, 116, 117, 59, 119] and Bayesian graph neural networks (BGNNs) [60, 120, 121, 61, 62, 63, 44, 133].
GVAEs. VGAE [58] is one of the earliest works to apply VAEs to graph data. In VGAE, a GCN-based encoder processes the input graph $G=(V,E)$ with node features $\mathbf{X}$ and adjacency matrix $\mathbf{A}$, and maps each node $i\in V$ to a latent variable $\mathbf{z}_{i}$. Instead of producing a fixed embedding, the encoder defines a probabilistic representation by estimating a Gaussian distribution for each node:
$q_{\boldsymbol{\phi}}(\mathbf{z}_{i}\mid\mathbf{X},\mathbf{A})=\mathcal{N}\!\bigl{(}\mathbf{z}_{i}\mid\boldsymbol{\mu}_{i},\,\mathrm{diag}(\boldsymbol{\sigma}_{i}^{2})\bigr{)},$ (16)
where the mean vector $\boldsymbol{\mu}_{i}$ represents the central embedding and the variance $\boldsymbol{\sigma}_{i}^{2}$ quantifies the uncertainty associated with node $i$’s representation. On the decoder side, an inner product decoder is typically employed to reconstruct the graph structure from these uncertainty-aware latent representations.
Although VGAE introduced probabilistic latent representations for graph, it still faces challenges in capturing complex dependency patterns, providing interpretable latent variables, and generating meaningful embeddings for isolated nodes. Consequently, various extensions have been proposed. Hasanzadeh et al. [114] propose a hierarchical variational framework that enables neighboring nodes to share information, thereby improving the modeling of the graph’s dependency structure. It allows the propagation of graph structural information and distribution uncertainty, which are essential for capturing complex posterior distributions. To clearly explain what latent factors are and why they perform well, Li et al. [115] introduces the Dirichlet Graph Variational Autoencoder (DGVAE) with graph cluster memberships as latent factors. They connect VAE-based graph generation with balanced graph cut, providing a new perspective on how to enhance the internal mechanisms of VAEs for graph generation. To avoid near-zero embeddings for isolated nodes generated by existing autoencoders, Variational Graph Normalized AutoEncoder (VGNAE) [116] utilizes $L_{2}$-normalization to derive improved embeddings for isolated nodes. Grover et al. [117] propose an “iterative” GNN-based decoder to progressively refine the reconstruction of graph adjacency matrices via low-rank approximations, enabling efficient and scalable unsupervised representation learning on large-scale graphs. Unlike the node-level methods described above, GraphVAE [59] modifies the encoder and decoder to work with latent representations of entire graphs. This approach focuses on modeling the uncertainty in graph-level representations, rather than in node-level ones.
Capitalizing on the capabilities of GVAEs in representation and uncertainty modeling, several studies have extended them to model dynamic graphs more effectively. Hajiramezanali et al. [118] develop a novel hierarchical variational model that introduces additional latent random variables. This model jointly captures the hidden states of a Graph Recurrent Neural Network (GRNN) to effectively represent both topological changes and node attribute variations in dynamic graphs. To learn dynamic graph representations in hyperbolic space while also modeling uncertainty, TGNN [119] incorporating a hyperbolic graph variational autoencoder is proposed, which encodes stochastic nodes from hyperbolic normal distributions.
BGNNs. A series of BGNN works consider model parameters, passed messages, and other key elements in graph learning as distributions rather than deterministic values. Bayesian GCNN [60] treats the observed graph $G_{obs}$ as a sample from an assortative mixed membership stochastic block model (a-MMSBM)  [134, 135]. As a result, the posterior probability of node (or graph) labels can be computed by:
$p(\mathbf{Y}^{*}|\mathbf{Y}^{*}_{\mathcal{L}},\mathbf{X},G_{obs})=$ $\int p(\mathbf{Z}|\boldsymbol{\theta},G,\mathbf{X})p(\boldsymbol{\theta}|\mathbf{Y}_{\mathcal{L}},G,\mathbf{X})$
$\cdot p(G|\lambda)p(\lambda|G_{obs})\,d\boldsymbol{\theta}\,dG\,d\lambda\,.$ (17)
Here $\boldsymbol{\theta}$ is a random variable representing the parameters of a Bayesian GCN over graph $G$, and $\lambda$ denotes the parameters that characterize a family of random graphs. $\mathbf{Y}_{\mathcal{L}}$ and $\mathbf{X}$ represent the training label set and the node feature matrix, respectively. $p(\mathbf{Y}^{*}|\boldsymbol{\theta},G,\mathbf{X})$ can be modeled using a categorical distribution by applying a softmax function to the output of the $L$-layer GCN $\mathbf{H}^{(L)}$. As an alternative to a-MMSBM, BGCN (node copying) [120] uses a node copying-based generative model, while BGCN (non-parametric) [121] employs a non-parametric generative model to construct the posterior distribution of graph adjacency matrices.
In addition to treating the observed graph as a random variable, the following works also assign probability distributions to other GNN components, adopting a Bayesian approach to model uncertainty.
To model the propagation of uncertainty in message-passing mechanisms, BUP [61] treats messages as multivariate Gaussian variables and employs GNNs to predict their means. Based on the assumption that additional neighboring nodes provide more evidence and reduce uncertainty, BUP defines an uncertainty propagation mechanism in GNNs for predicting the covariance of Gaussian distributions, which is different from traditional message passing. The covariance ultimately provides the uncertainty of predictions. GKDE [62] views model parameters as random variables and then obtains both EU and AU based on the uncertainty decomposition method in  [136, 137]. Additionally, GKDE incorporates uncertainty from evidence theory, namely vacuity resulting from lack of evidence and dissonance arising from conflicting evidence [138], offering a richer representation of uncertainty. [63] considered the adjacency matrix $\mathbf{A}$ as a random variable with prior distribution:
$p(\mathbf{A})=\prod_{ij}p(A_{ij}),p(A_{ij})=\text{Bern}(A_{ij}|\rho_{ij}^{o}),$ (18)
where $\text{Bern}(A_{ij}|\rho_{ij}^{o})$ is a Bernoulli distribution over $\rho_{ij}^{o}$. The variational posterior is given by:
$q_{\phi}(\mathbf{A})=\prod_{ij}q_{\phi}(A_{ij}),q_{\phi}(A_{ij})=\text{Bern}(A_{ij}|\rho_{ij}),\,\rho_{ij}>0,$ (19)
where $\rho_{ij}$ are free parameters, and $\phi=\{\rho_{ij}\}_{i,j}$.
To separately quantify aleatoric and epistemic uncertainty, BGNN [44] treats probabilistic links, noise in node features and model parameters as random variables. Aleatoric uncertainty is propagated through mechanisms defining the means and variances of node embeddings, while epistemic uncertainty is estimated using MC dropout. In contrast to BGNN, Blei et al. [133] propose to decompose the total uncertainty based on the Bayesian inference of node labels. Total uncertainty is the posterior probability over the remaining unobserved node labels conditioned on acquiring the target node’s label; aleatoric uncertainty is the same posterior before label acquisition; and epistemic uncertainty is their ratio, reflecting the information gain from observing that label. In this decomposition, they find that leveraging epistemic uncertainty for node selection more effectively guides graph active learning.
IV-B Gaussian Process
IV-B1 Gaussian Process Overview
Gaussian Processes (GPs) are a popular class of non-parametric Bayesian models that have been widely used for uncertainty quantification in machine learning [139]. A GP is defined as a collection of random variables, any finite number of which have a joint Gaussian distribution. In the context of function approximation, a GP can be used as a prior distribution over the space of functions, denoted as:
$f(\mathbf{x})\sim\mathcal{GP}(m(\mathbf{x}),k(\mathbf{x},\mathbf{x^{\prime}})),$ (20)
where the mean function $m(\mathbf{x})$ calculates the expectation of the function at input $\mathbf{x}$. The covariance function $k(\mathbf{x},\mathbf{x^{\prime}})$ captures the similarity between pairs of inputs and determines the smoothness of the functions sampled from the GP.
Given a set of observed data $\mathcal{D}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{n}$, where $\mathbf{x}_{i}$ are the input points and $y_{i}$ are the corresponding target values, the posterior distribution is a GP with updated mean and covariance functions:
$\mu(\mathbf{x})$ $=m(\mathbf{x})+\mathbf{k}(\mathbf{x})^{T}(\mathbf{K}+\sigma^{2}_{n}\mathbf{I})^{-1}(\mathbf{y}-\mathbf{m})$ (21)
$\Sigma(\mathbf{x},\mathbf{x^{\prime}})$ $=k(\mathbf{x},\mathbf{x^{\prime}})-\mathbf{k}(\mathbf{x})^{T}(\mathbf{K}+\sigma^{2}_{n}\mathbf{I})^{-1}\mathbf{k}(\mathbf{x^{\prime}})$ (22)
where $\mathbf{k}(\mathbf{x})=[k(\mathbf{x},\mathbf{x}_{1}),\dots,k(\mathbf{x},\mathbf{x}_{n})]^{T}$, $\mathbf{K}$ is the covariance matrix (or kernel matrix) with entries $\mathbf{K}_{ij}=k(\mathbf{x}_{i},\mathbf{x}_{j})$, $\mathbf{y}=[y_{1},\dots,y_{n}]^{T}$, $\mathbf{m}=[m(\mathbf{x}_{1}),\dots,m(\mathbf{x}_{n})]^{T}$, and $\sigma^{2}_{n}$ is the noise variance. The posterior distribution allows for uncertainty quantification in predictions. For a new input point $\mathbf{x}^{*}$, the predictive distribution is given by:
$p(f^{*}|\mathbf{x}^{*},\mathcal{D})=\mathcal{N}(f^{*}|\mu(\mathbf{x}^{*}),\Sigma(\mathbf{x}^{*},\mathbf{x}^{*})).$ (23)
The predictive mean $\mu(\mathbf{x}^{*})$ provides the expected value of the function at $\mathbf{x}^{*}$, while the predictive variance $\Sigma(\mathbf{x}^{*},\mathbf{x}^{*})$ quantifies the uncertainty associated with the prediction.
To address the computational cost of standard GPs, efficient approximations have been developed to preserve Bayesian uncertainty quantification at reduced cost [140, 107, 141]. Recent advancements have also explored the integration of GPs with deep learning architectures to leverage the strengths of both approaches [142, 143]. This integration enhances the ability to quantify uncertainty in complex data representations.
IV-B2 Gaussian Process for Graphs
Applying GPs to graph-structured data has emerged as an important research direction in recent years [64, 65, 122, 123, 124, 125, 126, 66, 127, 128]. This approach combines the powerful UQ capabilities of GPs with the complex structures of graph data, offering new possibilities for uncertainty estimation in graph learning tasks. The following discusses works closely related to UQ in graph learning, along with other relevant GP methods.
[64] were among the first to propose a Graph Gaussian Process (GGP) based semi-supervised learning method. This method can achieve good generalization performance even with limited labeled data, bridging the gap between simple models and data-intensive graph neural networks. GGP defines a GP prior over node embeddings, using a kernel function that can be selected from existing well-studied options. The model combines the GP prior with a robust-max likelihood [144] to handle multi-class problems. As an extension of the GGP model, UaGGP [65] addresses the uncertainty in graph structures that may arise in semi-supervised learning problems on graphs, where edges can connect data points with different labels, leading to uncertainty in predictions. UaGGP uses GP to guide the model learning by incorporating prediction uncertainty and label smoothness regularization. For any two nodes $v_{i}$ and $v_{j}$, UaGGP defines an uncertainty-aware distance measure $d(v_{i},v_{j})$ to constrain the proximity of the predictions at those nodes. This mechanism manifests uncertainty quantification by ensuring that $d(v_{i},v_{j})$ increases whenever the model’s confidence in one or both node predictions is low.
Subsequent extensions address scalability and specialized tasks: DGPG [125] proposes stochastic deep GPs for scalable node regression, while LNK [66] combines GNN-based encoders with sparse variational GPs for uncertainty estimation in molecular energy and force predictions. FT-GP and WT-GP [127] employ Fourier and wavelet transforms to capture multiscale and localized spectral information for graph classification. Similarly, GPG [124] integrates the graph Laplacian into the GP covariance matrix to enhance spectral regularization of target signals. GGPN [128] uses GPs to learn distributed embedding functions, enhancing flexibility in representing diverse graph structures.
IV-C OOD Epistemic Uncertainty Representation
Enhancing the generalization is an effective strategy for improving models’ performance under uncertainty. It enables models to make accurate predictions for “unexpected” and previously unseen samples. These outlier samples, which significantly differ from the training (in-distribution, ID) data, are commonly referred to as out-of-distribution (OOD) samples. In this subsection, we survey various approaches to model perturbations and represent outliers.
IV-C1 Dataset-level uncertainty estimation
One prominent approach to uncertainty representation is through dataset-level uncertainty estimation [145], where the model accounts for distributional shifts by defining an uncertainty set over data distributions. In this way, OOD or unseen data are modeled as elements outside a specified neighborhood of the empirical training distribution.
Formally, the uncertainty set is defined as $\mathcal{P}:=\{P|D(P,P_{0})\leq\rho\}$, where $P_{0}$ is the observed data distribution, and $\rho$ denotes a radius distance. The distance between two distributions, $P$ and $P_{0}$, is measured by a divergence metric $D(\cdot,\cdot)$. We list the most widely used choices of $D$.
[…]
  • KL-divergence, one special case of $f$-divergence, is used as a divergence measure [150, 67, 151, 152]. However, the possible limitations of metrics based on $f$-divergence, including KL-divergence, are that they may not be comprehensive enough to include some relevant distributions, or they may fail to precisely measure the divergence for extreme distributions [153].
[…]
  • Besides that, authors in [160] study uncertainty sets defined by maximum mean discrepancy, and Erdoğan and Garud [161] define the uncertainty sets using the Prohorov metric. Formally, $D_{Pro}(P,P_{0}):=\inf{\{\epsilon>0|P(X)\leq P_{0}(X^{\epsilon})+\epsilon\text{, }P_{0}(X)\leq P(X^{\epsilon})+\epsilon\}}$, where $X^{\epsilon}$ is the $\epsilon$-neighborhood of $X$.
  • Moreover, Wu et al. [162] define the graph-level distance by a novel graph subtree discrepancy based on Weisfeiler-Lehman Subtree Kernel. Fang et al. [152] define a graph-level heterophily distribution.
These dataset-level uncertainty estimation provides a principle way to incorporate distributional robustness into learning objectives, enabling models to represent and generalize better under real-world data variability and shift.
IV-C2 Sample-level uncertainty representation
Sample-level uncertainty refers to the uncertainty in the model’s prediction for one sample. For example, Qian et al. [145] measure it by perturbing an input sample and observing the variance in its predicted probabilities. This type of uncertainty quantifies how much the model’s prediction can change when the input is slightly modified. Adversarial samples are a prime illustration of such uncertainty.
By adopting adversarial attack methods such as FGSM [163], PGD [164], or CW attacks [165], one can generate adversarial samples that manipulate the model’s predictions by very limited perturbations. These adversarial samples essentially highlight regions where the model is uncertain and highly sensitive to small changes, and thus serve as a practical proxy to quantify sample-level uncertainty, with regions of the input space that produce significant shifts in model predictions being areas of high uncertainty. Formally, adversarial attacks can be defined in two alternative ways:
$\epsilon^{\ast}=\arg\max_{\|\epsilon\|\leq\rho}\mathcal{L}_{0}(\theta,x+\epsilon,y),$ (24)
or
$\epsilon^{\ast}=\arg\min_{\epsilon}\|\epsilon\|\;\;\textnormal{s.t. }f(x+\epsilon)\neq f(x),$ (25)
where $(x,y)$ denotes a sample with its corresponding label, and the sample $x$ is manipulated by an adversarial perturbation $\epsilon$, constrained by a predefined small budget $\rho$. $x+\epsilon^{\ast}$ is the adversarial counterpart of original input $x$. The general loss function $\mathcal{L}_{0}$, such as cross-entropy loss, quantifies the model’s performance on the specific task.
IV-C3 Manually designed uncertainty sets
OOD epistemic uncertainty arises from the lack of knowledge about certain inputs. When imposing a formal probability distribution on inputs or model parameters is hard, we can approximate this uncertainty by generating controlled perturbations of each input. These perturbations are carefully designed to preserve the core semantic or structural content while introducing realistic variations. By observing how the model’s predictions vary across these perturbations, we obtain a proxy measure of epistemic uncertainty.
Specifically, uncertainty sets are constructed by applying controlled perturbations while retain key structural or attributional information. By adopting controlled perturbations, input graphs are augmented in a manually controllable way. Structural augmentations include node/edge deletion or addition [91], and subgraph sampling via random walks to preserve local topology [74]. At the attribute level, node and edge features can be masked or corrupted to mimic missing or noisy information. Some works combine both structural and attribute perturbations to generate diverse graph augmentations [93].
These manually designed perturbations offer interpretable and task-relevant ways to simulate uncertainty, particularly when distributional assumptions are hard to formalize.
IV-D Stochastic Differential Equations
Stochastic Differential Equations (SDEs) offer a dynamic systems perspective for modeling uncertainty in neural networks. Instead of viewing network transformations as deterministic functions, SDE-based approaches model the evolution of hidden states or network parameters as stochastic processes. This continuous-time formulation provides a natural way to incorporate and quantify uncertainty throughout the network.
IV-D1 SDEs Overview
The core idea behind using SDEs for UQ in neural networks is to represent the mapping from input to output, or the learning dynamics of parameters, as the solution to an SDE. A general form of an SDE for a state process $\mathbf{x}_{t}$ is:
$d\mathbf{x}_{t}=f(\mathbf{x}_{t},t)dt+g(\mathbf{x}_{t},t)dW_{t}$ (26)
Here, $f(\mathbf{x}_{t},t)$ is the drift term, dictating the deterministic part of the evolution, and $g(\mathbf{x}_{t},t)$ is the diffusion term, scaling the stochastic part driven by a Wiener process (Brownian motion) $dW_{t}$. Both drift and diffusion functions can be parameterized by neural networks.
SDE-Net [166] instantiates the general SDE in Eq. (26) by using two neural networks. The first network (“drift net”) models the deterministic term. The second network (“diffusion net”) governs the stochastic term. The diffusion net induces a distribution over the final hidden state $\mathbf{x}_{T}$. Its variance, $\mathrm{Var}(\mathbf{x}_{T})$, quantifies epistemic uncertainty. Aleatoric uncertainty is then obtained from the predictive distribution $p(y\mid\mathbf{x}_{T})$, measured via expected predictive entropy for classification and predictive variance for regression.
Building on similar principles, other works have explored SDEs for UQ in various contexts. For instance, SDEs have been employed to conceptualize infinitely deep Bayesian neural networks, allowing for scalable approximate inference by leveraging a continuous-depth perspective [167]. Another line of work formulates UQ within stochastic optimal control problems, utilizing backward SDEs for gradient computation and uncertainty estimation [168]. Further research has focused on developing frameworks to explicitly quantify and propagate both aleatoric and epistemic uncertainties, such as using SDE-Recurrent Neural Networks for time-series data [169].
IV-D2 SDEs for Graphs
SDEs are increasingly employed with GNNs for UQ by modeling the evolution of graph representations as continuous-time stochastic processes. Bergna et al. propose the Latent Graph Neural Stochastic Differential Equations (LGNSDEs) framework to learn uncertainty-aware representations for graph data [76]. The core idea is to model the evolution of latent node features $\mathbf{Z}_{t}$ using an SDE. The framework defines an Ornstein-Uhlenbeck prior process and an approximate posterior SDE. The drift function $F_{G}$ of this posterior SDE is parameterized by a GCN, while the diffusion function $G_{G}$ is shared with the prior to ensure a bounded KL divergence. The approximate posterior SDE is given by:
$d\mathbf{Z}_{t}=F_{G}(\mathbf{Z}_{t},t,\boldsymbol{\phi})dt+G_{G}(\mathbf{Z}_{t},t)d\mathbf{W}_{t}$ (27)
where $\boldsymbol{\phi}$ represents the learned GCN weights and $d\mathbf{W}_{t}$ is a Wiener process.
GNSD [77] connects GNNs with Stochastic Partial Differential Equations (SPDEs). GNSD models an SPDE that ties the evolution of node representations to the graph structure. Drift and diffusion terms use graph operators such as Laplacians or neighborhood attention to drive node representation dynamics. For epistemic uncertainty, GNSD employs a specific stochastic forcing network and a Q-Wiener process defined on the graph domain to model its propagation among nodes, differing from LGNSDE’s Bayesian treatment of latent SDEs. In dynamic graph and generative modeling contexts, SDEs are utilized in decoders for inferring dynamic causal graphs in spatio-temporal forecasting [78] or to define continuous-time generative diffusion processes for permutation-invariant graph generation [79]. Furthermore, AGGDN model [80] stacks an SDE module upon an ODE component, with the ODE acting as a control signal to modulate the SDE propagation.
V Methods for Uncertainty Handling
Effective handling is essential for developing reliable machine learning models. In this section, we explore four key approaches for managing uncertainty: direct inference, out-of-distribution techniques, calibration, and conformal prediction. These methods offer distinct strategies for addressing scenarios where models directly handle uncertainty, encounter data outside its training distribution, adjusting model confidence, and ensuring robust predictions.
V-A Bayesian Inference Techniques
In this section we introduce methods for uncertainty handling based on Bayesian inference. We first discuss exact Bayesian inference methods that compute posterior distributions directly from Bayes theorem. We then present approximate methods that use Monte Carlo dropout and Variational Inference for graph data.
V-A1 Exact Bayesian Inference
Exact Bayesian Inference Overview. Exact Bayesian inference refers to methods that derive closed-form posterior distributions by directly applying Eq. (14) to the parameters of interest. This approach naturally fits node classification tasks, as neighboring nodes can be assumed as observed data (or evidence) to infer a node’s posterior distribution of the parameters used for classification.
Exact Bayesian Inference for Graphs. SocNL [81] is one of the early works based on exact Bayesian inference. It modeled node labels as categorical random variables and used a Dirichlet prior to initialize label distributions. The method assumes that a node tends to share similar labels with its neighbors, thus using the neighbors’ labels as observed data to update the prior and derive the closed-form Dirichlet posterior.
Beyond the label-sharing assumption and the mere propagation of labels (i.e., directly using the neighbors’ labels as observed data), NETCONF [83] introduces the propagation of multinomial messages to support both homophily and heterophily network effects. It captures these effects by employing compatibility matrices to represent the strength of connections between nodes and then using the resulting counts as evidence in the closed-form update. To handle high-dimensional node features, GPN [82] first uses an MLP to encode each node’s features. Then, normalizing flows [170] are applied to compute class pseudo-counts as observed data. Finally, Personalized PageRank propagates these class pseudo-counts to update the Dirichlet prior parameters for each node. SocNL, GPN, and NETCONF all utilize Dirichlet priors and posteriors, with the key distinction lies in the choice of observed data.
V-A2 Approximate Bayesian Inference
Approximate Bayesian Inference Overview. Approximate Bayesian inference is a well-established tool for handling uncertainty when exact Bayesian methods are computationally intractable. In this section, we focus on two prominent techniques: Monte Carlo (MC) Dropout [171] and Variational Inference (VI) [172].
MC Dropout approximates the Bayesian posterior by interpreting the dropout mechanism as a variational approximation over the model weights. Specifically, given an input $\mathbf{x}$, the predictive distribution is estimated by averaging over multiple stochastic forward passes as follows:
$p(\mathbf{y}\mid\mathbf{x})\approx\frac{1}{T}\sum_{t=1}^{T}p\bigl{(}\mathbf{y}\mid\mathbf{x},\hat{\boldsymbol{\theta}}^{(t)}\bigr{)},$ (28)
where $\hat{\boldsymbol{\theta}}^{(t)}$ represents a sampled set of model weights induced by applying dropout at each forward pass, $T$ is the total number of Monte Carlo samples.
In contrast, VI seeks to approximate the true posterior $p(\boldsymbol{\theta}\mid\mathcal{D})$ by a tractable variational distribution $q_{\boldsymbol{\phi}}(\boldsymbol{\theta})$, parameterized by $\boldsymbol{\phi}$. This approximation is achieved by maximizing the Evidence Lower BOund (ELBO):
$\mathcal{L}(\boldsymbol{\phi})=\mathbb{E}_{q_{\boldsymbol{\phi}}(\boldsymbol{\theta})}\Bigl{[}\log p(\mathcal{D}\mid\boldsymbol{\theta})\Bigr{]}-\mathrm{KL}\Bigl{(}q_{\boldsymbol{\phi}}(\boldsymbol{\theta})\parallel p(\boldsymbol{\theta})\Bigr{)},$ (29)
which balances the data likelihood with a regularization term expressed as the Kullback–Leibler divergence between the variational distribution and the prior $p(\boldsymbol{\theta})$. VI is also a cornerstone in approaches such as VAEs [56], where latent variable modeling plays a central role.
Approximate Bayesian Inference for Graphs. Graph learning models often employ approximate Bayesian inference techniques to handle uncertainty over model parameters, latent representations, and even the graph structure itself. In many cases, these techniques follow the same principles as those outlined in Eqs. (28) and (29), but they are adapted to account for graph data.
Bayesian GCNNs [60] leverage MC dropout to approximate the posterior distribution given in Eq. (IV-A2). The predictive posterior of node is estimated as
$p(\mathbf{Y}^{*}\mid\mathbf{Y}_{\mathcal{L}},\mathbf{X},G_{obs})\approx\frac{1}{N_{G}S}\sum_{i=1}^{N_{G}}\sum_{s=1}^{S}p\bigl{(}\mathbf{Y}^{*}\mid\hat{\boldsymbol{\theta}}_{s,i},G_{i},\mathbf{X}\bigr{)},$ (30)
where $G_{i}$ denotes a graph sample drawn from an assortative mixed membership stochastic block model (e.g., an a-MMSBM [134, 135]), and $\hat{\boldsymbol{\theta}}_{s,i}$ is a stochastic sample of weights (obtained via MC dropout). Similarly, other works [120, 121, 62, 44, 76] also use MC dropout to approximate the predictive posterior in their graph learning models.
VI is also applied in graph settings to approximate the posterior over diverse latent variables. For instance, VGAE [58] obtains the solution for Eq. (16) by optimizing the ELBO:
$\mathbb{E}_{q_{\boldsymbol{\phi}}(\mathbf{Z}|\mathbf{X},\mathbf{A})}[\log\,p(\mathbf{A}|\mathbf{Z})]-\text{KL}\Bigl{[}q_{\boldsymbol{\phi}}(\mathbf{Z}|\mathbf{X},\mathbf{A})\,\|\,p(\mathbf{Z})\Bigr{]}.$ (31)
$\log p(\mathbf{A}|\mathbf{Z})$ represents the reconstruction term, which encourages the model to accurately reconstruct the graph’s adjacency matrix. The KL divergence term $\mathrm{KL}[q(\mathbf{Z}|\mathbf{X},\mathbf{A})\|p(\mathbf{Z})]$ acts as a regularizer, ensuring that the learned distribution does not deviate too far from the prior $p(\mathbf{Z})$. Other works that similarly use VI to estimate the predictive posterior include approaches for node representations [114, 115, 116, 117, 118, 119, 64, 65, 66, 123, 125, 126, 127, 76, 78], for graph representations [59], and for graph structure [63].
V-B Handling uncertainty due to OOD data
When a model encounters OOD data, its epistemic uncertainty on those data points is typically higher. To handle the uncertainty, a surge of research has focused on OOD generalization, and most of them propose novel network architectures and objective functions to improve the model performance on OOD test data. We organize the discussion and introduce the studies according to the taxonomy of OOD handling strategies [173], which categories the methods into distributionally robust optimization (DRO) [146, 147, 148, 149, 150, 67, 153, 154, 155, 156, 160, 161, 174, 84, 68, 85, 69], adversarial training (AT) [86, 87, 175, 71, 176, 70, 88, 72, 89, 90], and self-supervised learning (SSL) [177, 178, 179, 180, 181, 182, 183, 184, 185, 186, 91, 74, 92, 75, 93, 187, 188].
V-B1 Distributionally Robust Optimization
Distributionally Robust Optimization Overview. Distributionally Robust Optimization (DRO) aims to enhance the model’s performance across a wide range of potential testing data distributions, instead of the training data distribution only. Specifically, DRO optimizes the objective function in the “worst case” of sample distributions $P\in\mathcal{P}$, in which place will the loss be maximized. Eq. (32) highlights the objective of DRO, which is formulated as a bi-level optimization problem:
$\min_{\boldsymbol{\theta}}\max_{P\in\mathcal{P}}\mathbb{E}_{(x,y)\sim P}[\mathcal{L}_{0}(\boldsymbol{\theta},x,y)],$ (32)
where $\mathcal{L}_{0}$ is a loss function. The uncertainty set $\mathcal{P}$, formally defined as $\mathcal{P}:=\{P|D(P,P_{0})\leq\rho\}$, encapsulates all the unseen data distribution within a specific distance from the training distribution. As discussed in Sec. IV-C1, several divergence metrics $D$ have been adopted in DRO.
Distributionally Robust Optimization for Graphs. Due to the unique topological structure of the graphs and various tasks involved, such as node classification and graph classification, the distributions could be defined on node and graph level, respectively. Specifically, the distance metric $D$ is employed to evaluate the discrepancy between distributions of node representations [84, 68, 85, 174], graph representations [67], or labels [69].
One widely used distance metric $D$ in DRO for graphs is the Wasserstein distance [68, 69, 85]. Specifically, Zhang, et al. [68] propose a graph learning framework comprising two models based on two separate assumptions regarding data uncertainty. One model assumes that every distribution within the uncertainty set follows a Gaussian distribution, while the other operates without any priors. In another work [174], the uncertainty set is defined in a Laplacian form. Besides, based on the assumption that nodes belonging to the same class follow the same underlying feature distribution, the Wasserstein distance is used to define the uncertainty set that includes the underlying distribution of node embeddings for each class [69].
Furthermore, some work [84, 67] explore the connection between DRO (using the KL-divergence as the distance metric) and graph learning methods. Specifically, Wu et al. [67] reveal that contrastive learning implicitly performs DRO over the negative sampling distribution. Then they refine the “worst-case distribution” and assign new weights to negative samples (in contrastive learning). In another work [84], LightGCN [^3] LightGCN only includes the neighborhood aggregation component, and excludes the feature transformation and nonlinear activation components. [189] is proven to work as a graph smoothness regularizer within the DRO framework, enabling a regularization-based method that promotes model robustness and generalization.
V-B2 Adversasrial Training
Adversarial Training Overview. Adversarial Training (AT) first generates adversarial samples, which aim to manipulate the original predictions and degrade model performance most with imperceptible perturbations. The models will then be trained with these adversarial samples. Formally, AT is formulated as the following min-max optimization problem:
$\min_{\theta}\max_{\|\epsilon\|\leq\rho}\mathcal{L}_{0}(\theta,x+\epsilon,y).$ (33)
Adversarial Training for Graphs. Considering both the attributional and topological features of graphs, AT for graphs considers the following min-max optimization problem:
$\min_{\boldsymbol{\theta}}\max_{\epsilon_{\theta},\epsilon_{x},\epsilon_{a}}\mathcal{L}_{0}(\boldsymbol{\theta}+\epsilon_{\theta},X+\epsilon_{x},A+\epsilon_{a},Y),$ (34)
where $\epsilon_{\theta}$, $\epsilon_{x}$, and $\epsilon_{a}$ represent the perturbations applied to model parameters $\boldsymbol{\theta}$, node feature matrix $X$, and adjacency matrix $A$, respectively. These perturbations are intended to be small enough and imperceptible, constrained by a budget such that $D(X+\epsilon_{x},X)\leq\rho_{x}$, with a slightly abuse of $D$. Similar constraints are applied to $\epsilon_{\theta}$ and $\epsilon_{a}$. Consequently, the uncertainties from model parameters, graph topologies, and node features are incorporated during the generation of adversarial samples. We summarize the existing studies from these three aspects.
A branch of AT studies on graphs focuses on manipulating model parameters $\boldsymbol{\theta}$ to improve generalization [86, 87], where the perturbations are evaluated by $L_{p}$ norm.
Some studies [175, 71, 176, 70, 88] focus on manipulating graph topologies, such as adding or deleting edges, corresponding to $\epsilon_{a}$ in Eq. (34). For the distance metric $D$, many works [175, 71, 176, 70] count the number of malicious edge modifications by the $L_{0}$ norm. In other work [88], discrete edge modifications are relaxed to continuous changes, evaluated using the $L_{1}$ norm.
Besides adding or deleting nodes from the graph, perturbing node features is a commonly explored strategy [72, 89, 90, 73], referring to $\epsilon_{x}$ in Eq. (34). Most relevant works [72, 89, 90] adopt the $L_{p}$ norm (e.g., $L_{1}$, $L_{2}$, or $L_{\infty}$) to constrain perturbations, and the Wasserstein distance has also been sparsely used [73].
Unlike DRO methods, which use distributional distance to set a boundary between certain and uncertain data, most AT studies rely on $L_{p}$ norm as distance metrics to capture potential variations in test data. While the $L_{p}$ norm could measure the magnitude of perturbations, it is not designed to quantify uncertainty. Instead, the $L_{p}$ norm defines a bounded region around an input, indicating the extent of perturbation, which implies the epistemic uncertainty of the model has about the data. Moreover, as previously mentioned studies [68, 69, 174], DRO frameworks leverage uncertainty sets to quantify epistemic uncertainty by measuring how much the model’s performance varies across these uncertainty sets. These approaches make it possible to model the knowledge gap the model has about unseen or shifted data distributions. However, most AT works focus on making models robust to data shift via adversarial perturbation, lacking direct uncertainty quantification studies by AT.
V-B3 Self-Supervised Learning
Self-Supervised Learning Overview. Self-Supervised Learning (SSL) allows models to learn informative signals from input samples and their variants, reducing the request for labeled samples, and thus alleviating the need for expensive annotations [177, 178]. SSL methods typically design pretext tasks that enable models to learn generalized representations from unlabeled data, providing a good initialization for downstream tasks. Various pretext tasks are proposed to extract meaningful features from data.
Taking SSL on computer vision as an example, the pretext tasks include image inpainting, where models learn to predict the missing parts masked by design [179]; image colonizing, where models learn to colorize grayscale images [180]; jigsaw puzzle solving, where models learn image representations by arranging scrambled image patches [181]; and self-supervised contrastive learning, which leverages image transformations to enhance feature learning [182, 183]. Besides, SSL methods have also seen remarkable success in the natural language processing domain [184, 185, 186].
Self-Supervised Learning for Graphs. SSL on graphs involves deliberately perturbing or modifying graphs to construct new controllable datasets while preserving the intrinsic information of the original graphs. These created variations of the graph instances allow SSL to simulate real-world shifts in data distributions. Thus, the model can better handle epistemic uncertainty by training on these augmented graphs. We categorize the SSL methods based on designing structural and attributional features.
Some studies focus on altering the structural features of graphs. For instance, You et al. [91] propose augmentations where a portion of nodes is randomly removed, and a set of edges can be either deleted or added. These modifications generate new graph instances while maintaining the majority of global structure. Additionally, they propose subgraph extraction using random walk, which is supposed to preserve the local structure of the original graph. The subgraph extraction augmentation method is also employed in [74] to capture local structure information effectively.
Other studies propose to manually design modification for the attributional features of graphs. Specifically, inspired by image inpainting, graph completion is proposed in [92], where the attributional features of nodes are masked, and the model learns to recover these missing features. Similarly, Hu et al. [75] propose randomly masking attributional features on nodes and/or edges to create variations in the graph without changing its overall structure. Furthermore, You et al. [92] propose node clustering and graph partitioning methods, where nodes are clustered based on feature similarity or topological connections, and cluster indices are used as pseudo-labels for SSL tasks. Some works, like [93], augment graph data at both the structural and attribute levels, providing a more comprehensive approach to generating new graph views.
The objectives of SSL on graphs can vary significantly. For example, Yehudai et al. [187] propose a novel SSL task centered on predicting the d-pattern tree descriptor, which captures specific topological patterns within the graph. Meanwhile, Liu et al. [188] address the distribution shift issue caused by high-confidence unlabeled nodes. They propose a collaborative framework based on information gain to reweight samples, thereby mitigating the effects of this shift. This work highlights how SSL can adapt to changes in data distribution and reduce uncertainty for OOD samples.
Like AT, SSL studies often incorporate useful augmentations for robustness. While these augmentations help simulate different data variations and implicitly introduce uncertainty, they do not explicitly measure or quantify the epistemic uncertainty caused by these transformations. Nonetheless, SSL augmentations are often more diverse and challenging due to significant perturbations over inputs. It allows SSL to better generalize models across a wider range of possible variations in graph data.
V-C Calibration
Calibration Overview.
Definition of Calibration. We begin with a classifier $f$ trained by a feature set and a corresponding label set whose label space is $\mathcal{Y}=\{1,...,K\}$. With inputs $x$ and the true label $y$, which can be represented as a one-hot vector $\textbf{y}=(y^{1},...,y^{K})$, $f$ applies softmax operation to output a probability vector, $\textbf{v}=(v^{1},...,v^{K}),\sum_{j=1}^{K}v^{j}=1$. The predicted label and confidence of $x$, denoted by $\hat{y}$ and $\hat{p}$, respectively, are
$\hat{y}=\arg\max_{j\in\mathcal{Y}}v^{j},\text{ }\hat{p}=\max_{j\in\mathcal{Y}}v^{j}.$ (35)
A model is considered well-calibrated if the confidence matches the probability of correct prediction [9]: $\text{Pr}(\hat{y}={y}|\hat{p}=p)=p,\forall p\in[0,1]$. A stronger definition of being well-calibrated requires that each class (not only the predicted label) [95, 190] are calibrated. More formally, $\text{Pr}(y=j|v^{j}=p)=p,\text{ }\forall j\in\mathcal{Y},\text{ }\forall p\in[0,1]$.
However, the calibration performance of modern neural networks has been proven poor along a series of factors, including network width, depth, batch normalization, and weight decay [9].
Calibration Methods. Calibration can be conducted on top of a trained accuracy-driven model to adjust prediction confidences via a calibration set $\{(x_{i},y_{i})\}_{i=1}^{n}$. In this case, calibration methods are called post-processing or post-hoc [94, 95, 191, 96, 192]. A popular non-parametric calibration approach within this group is Histogram Binning [94]. The idea is to group predictions into bins according to their confidence values. For instance, in a binary classification task with label space $\mathcal{Y}=\{0,1\}$, the probability vector of $x_{i}$ is $\textbf{v}_{i}=(v_{i}^{0},v_{i}^{1})$. Each bin $B_{m}$ is a set including samples whose $v_{i}^{1}$ is in the interval $I_{m}=(\frac{m-1}{M},\frac{m}{M}]$ based on a user-defined $M$. Therefore, we can use a calibration set to minimize
$\min_{\theta_{1},...,\theta_{M}}\sum_{m=1}^{M}\sum_{i\in B_{m}}(\theta_{m}-y_{i})^{2},$ (36)
where $\theta_{m}$ is the estimated number of positive-class samples in the $B_{m}$ set. For a test sample $x_{n+1}$ whose $v_{n+1}^{1}$ belong to $I_{m}$, $v_{n+1}^{1}$ should be replaced by $\theta_{m}$ as a calibrated confidence. Based on Histogram Binning, isotonic regression further optimizes the number and size of $I_{m}$ [95]. It is applied to deep learning models for regression tasks with promising performance [191]. Histogram Binning and isotonic regression can be applied to multi-class classification tasks using the one-vs-rest strategy. Nevertheless, both methods involve the creation of bins based on predicted confidence levels rather than directly on the input features, so the prediction’s confidence adjustment is determined by the bin to which it belongs, rather than by the characteristics of the input features themselves. As a result, neither method is adaptive to the input. In contrast, Platt scaling is a parametric input-adaptive calibration method [96, 193]. Traditionally, $\textbf{v}_{i}$ is transformed from a logit vector $\textbf{z}_{i}$ via softmax function $\sigma(\cdot)$. Platt scaling generalizes the process by $\textbf{v}_{i}=\sigma(\textbf{W}\cdot\textbf{z}_{i}+\textbf{b})$, which can be optimized by minimizing negative log-likelihood (NLL) on a calibration set. When W is a scalar and $\textbf{b}=0$, it is also referred to as temperature scaling. Accuracy versus uncertainty calibration (AvUC) loss [192] extends the temperature scaling method to encourage high confidence in accurate predictions and low confidence for likely incorrect predictions.
Some calibration techniques can be integrated directly into the training process, without requiring an additional calibration dataset [98, 97, 194]. With a training set $\{(x_{i},y_{i})\}_{i=1}^{n}$, a model is usually trained for a classification task NLL. However, NLL is minimized when the model assigns high probability to the correct class for each instance. In other words, the loss encourages the model to be confident in its predictions. While this is desirable when predictions are correct, it can lead to overconfident outputs even on uncertain or incorrect inputs. As a result, minimizing NLL too aggressively can lead to miscalibrated models that are overconfident and poorly represent uncertainty [194]. To mitigate the issue, label smoothing tunes a parameter $\alpha\in[0,1]$, so the predicted one-hot vector of $x_{i}$, denoted by $\textbf{y}_{i}=(y_{i}^{1},...,y_{i}^{K})$, becomes smoother by
$\mathbf{y}_{i}\rightarrow\mathbf{y^{\prime}}_{i}=\left((1-\alpha)y_{i}^{1}+\frac{\alpha}{K},...,(1-\alpha)y_{i}^{K}+\frac{\alpha}{K}\right),$ (37)
so it maps $\mathbf{y}_{i}$ to $\mathbf{y^{\prime}}$ by not assigning a full probability to the ground truth label [97]. Similarly, focal loss (FL) [98] is applied to calibration by
$\text{FL}(x_{i})=-(1-\hat{p}_{i})^{\gamma}\log(\hat{p}_{i}),\gamma\geq 0,$ (38)
where $\hat{p}_{i}$ is the prediction confidence of input $x_{i}$ [194]. Focal loss only considers confidence in the true label and the loss value will decrease faster as we increase $\gamma$, so it can prevent being over-confident. If $\gamma=0$, FL recovers CE.
Calibration for Graphs. Calibration on classification GNNs started to draw people’s attention recently as well [195, 99, 100, 101, 102, 130, 129, 103, 104, 37]. The works of calibrating GNNs can be categorized into node-level-classification, link prediction, and graph-level classification.
Node-level Classification. Traditional calibration techniques are usually applied to models handling independent and identically distributed (i.i.d.) data. In contrast, graph data presents a unique challenge due to its dependent nature. In graph data, the characteristics or status of one node may be influenced by the attributes or conditions of neighboring nodes, highlighting an inherent dependency where a node’s properties are not solely its own but also a reflection of its position within the graph [195]. This interdependency complicates the calibration process for GNNs. Furthermore, issues with class imbalance can lead to collapsed predictions, necessitating the adjustment of class weights during the training phase.
The calibration performance of GNNs on node-level classification tasks is related to GNNs’ architectures, neighbors’ labels, and the number of training epochs. Studies have shown that GNNs tend to produce underconfident predictions [99, 100]. Besides, it is proved that GNNs with wider layers tend to be better calibrated, and GNNs change from being under-confident to overconfident as the number of network layers increases [101, 102]. Calibration error in the classification of a node can be influenced by the node’s neighbor as well. Experimental results show that well-calibrated confidence distributions in graphs are often homophilic [99]—that is, neighboring nodes tend to have similar ground-truth confidences. Furthermore, it has been shown that calibration error decreases as the number of same-label neighbors increases [100], suggesting that strong local label agreement improves calibration. However, Liu et al. [101] report a contrasting observation: well-calibrated predictions are more likely when the same-class neighbor ratio is moderate (around 0.4 to 0.5), and models tend to become overconfident when the same-class neighbor ratio is too high. This apparent conflict between favoring high versus moderate same-label neighbor ratios may stem from differences in graph structures, model architectures, or training regimes (e.g., number of epochs). Besides, Liu et al. [101] show that GNNs tend to be under-confident with early stopping and will be overconfident with too many training epochs.
Most uncertainty quantification methods designed for multi-layer perceptrons can not recalibrate GNNs well. Teixeira et al.[195] and Vos et al.[196] show that methods like histogram binning, isotonic regression, Platt scaling, Accuracy-versus-Uncertainty (AvU)[197], and Graph Calibration Loss (GCL)[102]—which will be introduced shortly—generally outperform approaches such as Monte Carlo dropout [171] and model ensembles [198]. While Monte Carlo dropout and ensembles can improve calibration by capturing model uncertainty through stochastic regularization and diversity across multiple models, respectively, their effectiveness is often limited by computational cost and the lack of post-hoc adjustment. In contrast, the former methods directly refine the confidence estimates, leading to better-calibrated predictions. Moreover, models optimized on multi-class node classification tasks will be better calibrated and more discriminative than those trained on binary classification tasks [196]. Based on these insights, some works developed calibration approaches customized to GNNs.
The calibration methods for GNNs can be post hoc [99, 100, 101]. CaGCN is a topology-aware post-hoc calibration GCN model assuming that neighboring nodes share similar ground truth confidence levels with the center node [99]. With an undirected attributed GCN optimized for prediction accuracy, CaGCN works as another GCN to calibrate the outputs of the accuracy-driven model. Hence, CaGCN can be regarded as a Platt scaling method adapted to graph models. To preserve the accuracy of the original classification GCN, the optimization objective function of CaGCN is a combination of negative log-likelihood loss and a customized calibration loss, which aims to calibrate under-confident correct predictions and over-confident incorrect predictions. Graph Attention Temperature Scaling (GATS) [100] is also a post-hoc method, and embeds the attention mechanism for node-wise calibration. For each node, a scaling process will consider global calibration bias, relative confidence compared with neighbors, and aggregated neighborhood influence with corresponding attention coefficients. Ratio-binned scaling (RBS) follows the logic of histogram binning to calibrate outputs of GNNs [101]. RBS first estimates the same-class-neighbor ratio of each node, then groups the nodes into bins according to their ratio values, and finally learns a temperature for each bin on a validation set.
Conducting calibration during model training is also attractive since it does not require additional data and computation costs. Graph Calibration Loss (GCL) enables an end-to-end calibration method for GNNs [102]. GCL minimizes the discrepancy between the ground-truth and predicted confidence distributions. Specifically, GCL is the KL divergence between the two distributions regularized by the product of their entropies. HyperU-GCN is proposed to calibrate the hyperparameter uncertainty of automated GNNs using bi-level optimization [199]. The upper-level problem explains uncountable uncertainties by developing probabilistic hypernetworks through a variational Bayesian method. The lower-level problem learns how the hyperparameter distribution is propagated to GCN weights.
To maintain well-calibrated predictions in the presence of adversarial attacks and discriminatory biases, Multi-view Confidence-calibrated NIFTY (MCCNIFTY) was introduced [129]. Previous methods, such as FairGNN [200], RobustGNN [201], and NIFTY [202], primarily focused on preserving accuracy under such conditions, but often produced overconfident incorrect predictions or were vulnerable to global graph noise. MCCNIFTY combines intra-view calibration, inter-view fusion, and uncertainty-aware message passing to generate well-calibrated predictions under the impact of adversarial perturbations and biases.
Link Prediction. Link prediction is an important downstream task of GNNs [37]. Recent research finds that GNNs tend to be overconfident in negative link predictions but underconfident in positive ones [130].
IN-N-OUT [130] calibrates link prediction by learning a temperature parameter through edge perturbations. In this approach, the embedding of an edge in a GNN depends on the predicted labels of its two connected nodes. During training, two distinct embeddings are computed for each edge: one with the edge present in the graph (included in message passing) and one with the edge removed (excluded from message passing). The discrepancy between these two embeddings is utilized to determine the appropriate value for temperature scaling. For example, a high discrepancy between the two embeddings suggests that the confidence associated with the existence of the edge should be reduced.
In contrast, E-$\Delta$UQ [103] calibrates link prediction by perturbing node features. It is based on an anchoring technique, called $\Delta$-UQ [203]. Specifically, $\Delta$-UQ randomly chooses an anchor $c$, which is a constant bias with the same size of input feature, for stochastic centering during training, and creates a relative representation $[x-c,c]$ of each input $x$. Therefore, a model is trained on $([x-c,c],y)$ instead of $(x,y)$. The anchor value $c$ is different in each training iteration. During inference, $\Delta$-UQ generates multiple predictions for an input $x$ using different random perturbations $c$. If the model is confident about a particular label, the corresponding softmax scores remain stable across these random anchoring. In contrast, uncertain predictions exhibit higher variability. The final calibrated confidence distribution over labels is obtained by subtracting the standard deviation from the mean of the predicted softmax distributions across perturbations. The calibrated output of input $x$ is the label with the highest value in this adjusted confidence distribution. E-$\Delta$UQ extends the method to calibrate link prediction through two stages: encoding node information and decoding to predict links between node pairs. Random anchoring is applied during the encoding and/or decoding stages to enable uncertainty-aware training. It has three variants. The first variant implements anchoring between encoding and decoding, thus it is based on a deterministic encoder but the decoder is stochastic. The second variant implements one more anchoring after specific layers of encoding. Based on the second one, the last variant pre-trains the encoder with an auxiliary attribute masking task on node features. Although both IN-N-OUT and E-$\Delta$UQ rely on perturbation, E-$\Delta$UQ offers greater flexibility since its anchoring is not limited to binary perturbations as in IN-N-OUT.
Graph-level Classification. Under the conditions of covariate and concept shifts, the anchoring technique is also adopted in G-$\Delta$UQ [104] with three variants for the calibration of GNNs on graph-level classification using Superpixel-MNIST datasets [204, 205, 206, 207]. The first variant, node feature anchoring, applies stochastic centering directly to the input node features. The second, intermediate anchoring, introduces stochastic centering after specified hidden layers for each node. The final variant, readout anchoring, applies anchoring to the overall graph representation before the graph-level readout function, rather than to individual nodes.
V-D Conformal Prediction
Conformal Prediction Overview. Given a test input, conformal prediction (CP) generates a prediction set (rather than a single prediction) based on a statistical threshold derived from uncalibrated confidences of previous predictions. The prediction set guarantees the inclusion of the true label with a specified probability [105] or an upper bound of a user-defined prediction loss [208]. We concentrate on the use of CP in classification tasks, as GNNs are commonly employed in this context.
Inductive CP, or split conformal prediction (SCP), is a widely applied version of CP [106]. For a trained classifier ${f}$ and input $x$ with label space $\mathcal{Y}=\{1,...,K\}$, the estimated probability for each class is $\textbf{v}=(v^{1},...,v^{K})$. Let $f(x)_{y}=v^{y}$ denote the predicted probability of the true class $y$ from the softmax layer. The calibration set $D_{cal}=\{(X_{i},Y_{i})\}_{i=1}^{n}$ is applied to quantify how uncertain the model $f$ would be. The score function $s(x,y)=1-f(x)_{y}$ is widely selected for classification tasks, and it measures how predictions conform to the true values at inputs. The conformal score for $(X_{i},Y_{i})\in D_{cal}$ is defined as $s_{i}=1-f(X_{i})_{Y_{i}}$. Let $S=\{s_{i}\}_{i=1}^{n}$, we define $\tau$ as the empirical $1-\alpha$ quantile of $S$. In other words, $\tau=\lceil(1-\alpha)(n+1)\rceil$-th smallest element of $S$. Formally, defining $\delta_{s_{i}}$ as a point mass at $s_{i}$, the quantile is
$\tau=\text{Quantile}\left(1-\alpha,\frac{1}{n}\sum\nolimits_{i=1}^{n}\delta_{s_{i}}\right).$ (39)
For a test input $X_{n+1}$, a prediction set ${C}(X_{n+1})=\{k|1-f(X_{n+1})_{k}\leq\tau,k\in\mathcal{Y}\}$ includes all classes $k\in\mathcal{Y}=\{1,...,K\}$ whose conformal score is less or equal to $\tau$. If the test sample $(X_{n+1},Y_{n+1})$ is exchangeable with the elements of $D_{cal}$, a coverage guarantee holds that
$\text{Pr}\left(Y_{n+1}\in C(X_{n+1})\right)\geq 1-\alpha.$ (40)
Transductive CP, or full conformal prediction (FCP), is one of the earliest approaches to CP [105]. With training data $D_{train}=\{(X_{i},Y_{i})\}_{i=1}^{n}$ and a test input $X_{n+1}$, the trained model $f^{k}$ is based on $D_{train}\cup\{(X_{n+1},k)\},\forall k\in\{1,2,...,K\}$, so multiple models need to be trained. For each $f^{k}$, the conformal scores are $s_{i}^{k}=1-f^{k}(X_{i})_{Y_{i}}$ and $s_{n+1}^{k}=1-f^{k}(X_{n+1})_{k}$. If $s_{n+1}^{k}$ is less than the $1-\alpha$ quantile of $\{s_{i}^{k}\}_{i=1}^{n+1}$, we put the corresponding $k$ to the prediction set of $X_{n+1}$.
To make the size of $C(X_{n+1})$ more adaptive to $X_{n+1}$ under the exchangeability assumption, novel conformal score functions are proposed  [109, 209, 108, 210]. However, the exchangeability assumption can be violated by distribution shift between calibration distribution $P_{XY}$ and test distribution $Q_{XY}$ such that $(X_{n+1},Y_{n+1})$ is non-exchangeable with the elements of $D_{cal}$. In the case of covariate shift ($P_{X}\neq Q_{X}$), an importance weighting technique is introduced based on likelihood ratio to maintain the coverage guarantee in Eq. (40[211]. When joint distribution shift ($P_{X}\neq Q_{X}$, $P_{Y|X}\neq Q_{Y|X}$) occurs, prediction sets can be modified to keep $1-\alpha$ coverage on test data under dynamic distribution shift [212, 213], and static distribution shift [110, 214, 215]. It is proved that the coverage gap can be bounded by the total variation distance [216] and Wasserstein distance [217] between calibration and test conformal scores under the non-exchangeable condition.
Conformal Prediction for Graphs. When labeled calibration nodes and test nodes are randomly distributed throughout the graph, it is reasonable to assume that calibration and test nodes are exchangeable. Under this setting, marginal coverage guarantees remain valid. JuryGCN [111] estimates prediction uncertainty in node classification tasks using jackknife methods. In graph settings, node homophily plays an important role: if a node is surrounded by neighbors with similar labels, their features provide strong predictive cues, resulting in higher model confidence. Conversely, when a node is surrounded by neighbors with different labels (i.e., heterophily), the uncertainty in prediction should increase. Motivated by this observation, Diffused Adaptive Prediction Sets (DAPS) [112] leverage the graph structure via neighborhood diffusion and define the diffused score as:
$s^{\text{diff}}(X_{i},Y_{i})=(1-\beta)s(X_{i},Y_{i})+\frac{\beta}{|\mathcal{N}_{i}|}\sum{X_{j}\in\mathcal{N}_{i}}s(X_{j},Y_{i}),$
where $\beta$ signifies the diffusion parameter and $\mathcal{N}_{i}$ includes the features of node $i$ neighbors. The diffused score function $s^{\text{diff}}$ preserves exchangeability if the original scores $s(X_{i},Y_{i})$ are exchangeable. This diffusion approach is particularly beneficial in homophilic graphs, where neighboring nodes tend to share similar ground-truth labels. In such settings, the prediction for a node can be improved by incorporating the information from its structurally close, and often label-consistent, neighbors.
However, when calibration and test data correspond to different subgraphs, such as disjoint node sets with limited connectivity, the exchangeability assumption typically breaks down due to structural and distributional differences between subgraphs. This is because the calibration nodes and test nodes are no longer drawn from the same distributional context within the graph, especially if they differ in topological properties, feature distributions, or label characteristics. For example, if the calibration nodes all lie within a densely connected community or are concentrated in a homophilic region of the graph, while the test nodes are located in sparser or more heterophilic regions, the conformity scores computed on the calibration set may not generalize to the test nodes. This mismatch undermines the core assumption of exchangeability required for valid coverage guarantees. Furthermore, the violation of exchangeability is exacerbated in inductive settings where the test node influences (or is influenced by) calibration nodes through shared neighbors or overlapping receptive fields in message-passing architectures. This issue has been highlighted in recent works such as [112], which emphasize that topological locality and data partitioning within graphs introduce structured dependencies that vanilla CP methods are not designed to handle. Beyond the standard assumption of exchangeability, Barber et al. [216] propose a generalization of conformal prediction known as Non-exchangeable Conformal Prediction (NexCP), which accommodates settings where data points are not identically distributed. Instead of treating all calibration samples equally, NexCP assigns deterministic weights $w_{1},\dots,w_{n}\in[0,1]$ to $n$ calibration samples, and generates the prediction set as
${C}(X_{n+1})=\{k$ $|s(X_{n+1},k)\leq$ (41)
$\text{Quantile}\left(1-\alpha,\sum\nolimits_{i=1}^{n}w_{i}\delta_{s_{i}}\right),k\in\mathcal{Y}\}.$
The core challenge lies in selecting appropriate weights $w_{i}$ that reflect the similarity between calibration samples and the test input $X_{n+1}$. Intuitively, calibration examples that come from distributions closer to that of the test sample should receive higher weights, as they are more informative and relevant for estimating uncertainty. This idea has been extended to graph data by Clarkson et al.[113], who integrate NexCP into the Adaptive Prediction Sets (APS) framework [109]. In graphs, homophily is a common structural property. Leveraging this, Neighborhood Adaptive Prediction Sets (NAPS) are proposed to assign weights to calibration nodes based on their distances to the test node. Nodes that are closer (in terms of hop count or diffusion distance) are considered more exchangeable and are thus given greater influence in determining the prediction set. Three variants of NAPS have been developed, each capturing different notions of neighborhood similarity and diffusion dynamics within the graph. The primary variant, NAPS, assigns the weights $w_{i}=1$ if $X_{i}\in\mathcal{N}_{n+1}^{K}$, where $\mathcal{N}_{n+1}^{K}$ includes all nodes within $K$-hop neighborhood of node $X_{n+1}$. The second one, NAPS-H, adopts a hyperbolic decay rate for $w_{i}=k^{-1}$ where $k$ is the $k$-hop neighbor of node $X_{n+1}$. The third one, NAPS-G, uses a geometric decay rate for $w_{i}=2^{-k}$.
In addition to node classification, conformal prediction has been applied to other graph-related tasks, such as link prediction [218, 219, 220] and graph regression [221, 222]. However, these studies largely extend standard inductive CP to new task settings while still relying on the assumption of exchangeability. As a result, we do not elaborate on these works further.
VI Uncertainty Quantification Evaluation
This section discusses the evaluation of uncertainty quantification, including methods for direct inference and OOD data, calibration and conformal prediction.
VI-A Evaluating uncertainty due to OOD data
Unlike calibration and conformal prediction approaches that explicitly quantify and optimize predictive uncertainty, DRO, AT and SSL methods primarily target improved generalization under distributional shifts, e.g., OOD scenarios or adversarial attacks, without directly modeling uncertainty. In particular, DRO typically assesses model robustness by evaluating performance across shifted data distributions, while AT focuses on enhancing resilience to adversarial perturbations, evaluating both clean and adversarial predictive performance. Consequently, the evaluation of these techniques through the lens of uncertainty remains underexplored, highlighting a critical need for rigorous uncertainty-centric assessment frameworks.
VI-B Calibration
The evaluation of calibration measures the difference between a model’s prediction confidence and the exact probability that the predicted output is correct. Two widely applied metrics are introduced below.
Expected Calibration Error (ECE). The entire confidence space [0,1] can be divided into $M$ equal-size intervals $I_{m}=(\frac{m-1}{M},\frac{m}{M}]$ for $m=1,...,M$. $B_{m}$ is the set of samples with confidences in $I_{m}$. The accuracy and confidence of $B_{m}$ are
$\text{acc}(B_{m})=\frac{1}{|B_{m}|}\sum\nolimits_{i\in B_{m}}\textbf{1}(\hat{y}_{i}=y_{i}),$ (42)
$\text{conf}(B_{m})=\frac{1}{|B_{m}|}\sum\nolimits_{i\in B_{m}}\hat{p}_{i},$ (43)
where $y_{i}$ and $\hat{y}_{i}$ are the true and predicted label of input $x_{i}$, and $\hat{p}_{i}$ is the confidence of the prediction. Expected Calibration Error (ECE) computes the absolute difference between $\text{acc}(B_{m})$ and $\text{conf}(B_{m})$:
$\text{ECE}=\sum\nolimits_{m=1}^{M}\frac{|B_{m}|}{n}|\text{acc}(B_{m})-\text{conf}(B_{m})|.$ (44)
A reliability diagram can be plotted with $\text{acc}(B_{m})$ and $\text{conf}(B_{m})$ to visualize a model’s calibration performance [223]. The plot of a well-calibrated model should follow a diagonal, indicating the accuracy is the same as confidence. Some works extend the concept of ECE to make it suitable for different scenarios. For example, Maximum Calibration Error (MCE) computes the maximum gap between accuracy and confidence for $m=1,...,M$ [224], and Classwise ECE (cwECE) is designed to see averaged ECE on class level [225].
To evaluate the calibration of node-level classification models on GNNs, node-wise ECE follows the concept of ECE. Specifically, test nodes are partitioned into bins according to their prediction confidences, and nodewise-ECE works as Eq. (44) to quantify the summation of the absolute difference between confidence and accuracy across all bins [99, 102, 195, 101]. However, nodewise-ECE neglects graph structure and the interdependencies between nodes. To remedy this, edge-wise ECE is proposed in [226]. With graph $G=(\mathcal{V},\mathcal{E})$, denote $\mathcal{E}_{U}=\{(i,j)\in\mathcal{E}|i,j\in U\}$ the subset of edges $\mathcal{E}$ constrained to test nodes $U$. Given a trained model, the confidence of edge $(i,j)\in\mathcal{E}_{U}$ is $\hat{c}{(i,j)}=\hat{p}_{i}\hat{p}_{j}$. The calculation of edge-wise ECE is similar to node-wise ones. Specifically, each $\hat{c}{(i,j)}$ is computed and assigned to corresponding bins, and a prediction is considered accurate only if both predictions are correct (i.e. $y_{i}=\hat{y}_{i},y_{j}=\hat{y}_{j}$). The summed gap between confidence bins and accuracy is edge-wise ECE, similar to Eq. (44). Additionally, ECE for agreeing node pairs and disagreeing node pairs are also proposed [226].
Brier Score. As ECE depends on the resolution of the finite number of bins, a continuous proper scoring rule, such as Brier Score (BS) [227], can act as an alternative to ECE. BS considers a sample’s confidence in all labels. It is given by
$\text{BS}=\frac{1}{n}\sum_{i=1}^{n}\sum_{j=1}^{K}(v_{i}^{j}-\mathbbm{1}(y_{i}=j)),$ (45)
where $v_{i}^{j}$ is the probability that sample $i$ is labeled $j$ and $\mathbbm{1}(\cdot)$ is indicator function. A lower Brier Score indicates not only that a model is well-calibrated but also that the model is confident in its predictions.
VI-C Conformal Prediction
Conditional Coverage Metric. The key contribution of CP is the guaranteed coverage of prediction sets, as stated in Eq. (40). However, it is not explicitly claimed but desired that CP can be adaptive for different inputs (i.e. larger prediction sets for more uncertain predictions) [209]. In other words, the $1-\alpha$ coverage can be robust and stable no matter how the test domain changes. Let $E_{1},...,E_{N}$ denote the $N$ sets of test inputs coming from the corresponding test domains, and we want the coverage guarantee to always hold that $\forall e\in\{E_{1},...,E_{N}\}$:
$\text{Pr}(Y_{test}\in{C}(X_{test}))\geq 1-\alpha,X_{test}\in e.$ (46)
This test domain can be characterized by an explicit feature in input data or an underlying environment variable related to the data generation process. The worst coverage among multiple domains is given by
$\min_{e\in\{E_{1},...,E_{N}\}}\frac{1}{|\mathcal{I}_{e}|}\sum\nolimits_{i\in\mathcal{I}_{e}}\mathbbm{1}\{Y_{i}\in{C}(X_{i})\},$ (47)
where $\mathcal{I}_{e}$ represent the sample index set of set $e$ and $|\mathcal{I}_{e}|$ denote the number of samples in set $e$ [228, 229]. If the metric value is much lower than the expected coverage $1-\alpha$, the CP procedure performs poorly on conditional coverage.
Coverage Correctness. The empirical coverage by split conformal prediction involves randomness due to the limited number of available samples and possibly biased data splitting. Running conformal prediction multiple times can help determine the prediction coverage’s correctness and stability. With $n$ labeled samples available, we can randomly split them into $n_{cal}$ samples for calibration and $n_{val}$ samples for validation, and calculate the coverage by
$c_{t}=\frac{1}{n_{val}}\sum_{i=1}^{n_{val}}\mathbbm{1}\{Y_{i,t}\in{C}(X_{i,t})\},$ (48)
where $Y_{i,t}$ and $X_{i,t}$ are the label and feature of the i-th element of $n_{val}$ validation samples at $t$-th time splitting. After repeating for $T$ times, the final averaged empirical coverage, $\bar{c}$, will be $\frac{1}{T}\sum\nolimits_{t=1}^{T}c_{t}.$ The variance of these empirical coverage values indicates the stability as $\frac{1}{T}\sum_{t=1}^{T}(c_{t}-\bar{c})^{2}$.
VII Future Directions
In this section, we analyze existing challenges in uncertainty quantification on graphs, and summarize some future research directions facing the challenges.
Bayesian learning can generate credible regions for model predictions, but these regions are only valid if the model is correctly specified, meaning that the prior, likelihood, and computational capability must be accurately defined during posterior inference [230, 231, 232]. [230] primarily addresses these issues by introducing a temperature parameter into Bayesian GNNs within the conformal prediction framework. The combination of Bayesian inference with the conformal prediction framework allows for both credible regions and prediction sets that are more robust, even in the presence of model misspecification. Integrating Bayesian methods with other uncertainty quantification techniques could further mitigate inaccuracies arising from violations of Bayesian learning assumptions. By leveraging the strengths of complementary uncertainty quantification methods, such integrations could lead to more accurate and reliable uncertainty estimates for graph learning.
Out-of-distribution tasks have been well explored by various techniques, such as DRO, AT, and SSL. As discussed before, the majority of existing research has primarily focused on improving model generalization to OOD samples. However, there has been limited exploration into the deeper connections between UQ and OOD. A promising direction is to leverage OOD methods to better estimate and quantify uncertainties. Integrating OOD techniques with UQ could yield valuable insights, potentially leading to more robust and reliable uncertainty estimations. The integration could also help bridge the gap between OOD and UQ.
Conformal prediction for dynamic graphs presents a unique challenge. Most existing conformal prediction methods for graph data concentrate on static graphs within transductive and inductive frameworks to adhere to the exchangeability assumption or fixed distribution shift. However, dynamic graphs, such as those found in social networks, are common in real-world applications. These can result in dynamic distribution shifts, posing greater challenges for CP [233]. Addressing this challenge could enhance the interpretability of GNNs when dealing with evolving data, like changes in edges over time.
Calibration methods primarily address miscalibration errors by leveraging the homophily property of graphs, which assumes that neighboring nodes tend to belong to the same class [100, 99]. This approach has proven effective in many scenarios, but it falls short in heterophilous environments where neighboring nodes are more likely to belong to different classes. Despite the growing importance of heterophilous graphs, the calibration of GNNs in these contexts remains under-explored. Effectively calibrating GNNs under heterophily is essential for improving their reliability and performance in real-world applications characterized by complex relationships.
VIII Conclusion
In this survey, we carefully reviewed the uncertainty quantification within graphical models, specifically focusing on Graph Neural Networks and Probabilistic Graphical Models. We examined existing methods for representing and handling uncertainty, including Bayesian representation learning, Gaussian Processes, Outlier Representation, approaches based on Stochastic Differential Equations (SDEs), conformal prediction, calibration, and so on. The domain of uncertainty in graphical models presents substantial opportunities for further innovation and advancement. Advancing our understanding of uncertainty modeling in graph-based systems will empower machine learning systems to make more reliable decisions in the presence of real-world noise. Such progress will not only improve the accuracy and safety of AI applications but also promote trust and encourage the broader adoption of these technologies in practical settings.
Appendix A Resources
For convenient access, Table IV and Table V summarize open-source code repositories from some cited papers on related topics and tutorials from prominent machine learning conferences, respectively.
TABLE IV: Repositories of Open-source Codes
Method Model Algorithm Code Link
Bayesian Method General Model VAEs [56] https://github.com/AntixK/PyTorch-VAE
BNNs [131, 132] https://github.com/Harry24k/bayesian-neural-network-pytorch
Graphic Model NETCONF [83] https://dhivyaeswaran.github.io/code/netconf.zip
GPN [82] https://www.daml.in.tum.de/graph-postnet
VGAE [58] https://github.com/tkipf/gae
DGVAE [115] https://github.com/xiyou3368/DGVAE
SIG-VAE [114] https://github.com/sigvae/SIGraphVAE
VGNAE [116] https://github.com/SeongJinAhn/VGNAE
Graphite [117] https://github.com/ermongroup/graphite
VGRNN [118] https://github.com/VGraphRNN/VGRNN
Bayesian GCNN [60] https://github.com/huawei-noah/BGCN
Bayesian GCNN (node copying) [120] https://github.com/floregol/BGCN_copying
S-BGCN-T-K [62] https://github.com/zxj32/uncertainty-GNN
VGCN [63] https://github.com/ebonilla/VGCN
GGP [64] https://github.com/yincheng/GGP
GPG [124] https://www.kth.se/ise/research/reproducibleresearch1.433797
DGPG [125] https://github.com/naiqili/DGPG
FT-GP & WT-GP [126] https://github.com/FelixOpolka/Graph-Classification-Gaussian-Processes-via-Spectral-Features
LNK [66] https://github.com/wollschl/uncertainty_for_molecules
GGPN [128] https://github.com/sysu-gzchen/GGPN
Conformal Prediction General Model CPCS [211] https://github.com/ryantibs/conformal/tree/master/tibshirani2019
CQR [108] https://github.com/yromano/cqr
APS [109] https://github.com/msesia/arc
conformalbayes [234] https://github.com/CoryMcCartan/conformalbayes
EnbPI [213] https://github.com/hamrel-cxu/EnbPI
CP with Missing Values [235] https://github.com/mzaffran/ConformalPredictionMissingValues
LCP [210] https://github.com/LeyingGuan/LCP
stabCP [236] https://github.com/EugeneNdiaye/stable_conformal_prediction
OQR [229] https://github.com/Shai128/oqr
Graphic Model NAPS [113] https://github.com/jase-clarkson/graph_cp
DAPS [112] https://github.com/soroushzargar/DAPS
CF-GNN [237] https://github.com/snap-stanford/conformalized-gnn
JuryGCN [111] https://github.com/BlueWhaleZhou/JuryGCN_UQ
Calibration General Model Dirichlet Calibration [225] https://github.com/dirichletcal/dirichletcal.github.io
Calibrated Regression [191] https://github.com/AnthonyRentsch/calibrated_regression
Focal Calibration [194] https://github.com/torrvision/focal_calibration
Label Smoothing [97] https://github.com/seominseok0429/label-smoothing-visualization-pytorch
Spline Calibration [238] https://github.com/kartikgupta-at-anu/spline-calibration
Graphic Model G-$\Delta$UQ [104] https://github.com/pujacomputes/gduq
CaGCN [99] https://github.com/BUPT-GAMMA/CaGCN
GATS [100] https://github.com/hans66hsu/GATS
RBS [101] https://github.com/liu-yushan/calGNN
HyperU-GCN [199] https://github.com/xyang2316/HyperU-GCN
TABLE V: Tutorials on Uncertainty Quantification and Graphic Models
Tutorial $G$ Bayes CP Cal OOD Venues Year
Graph Learning: Principles, Challenges, and Open Directions ICML 2024
Formalizing Robustness in Neural Networks: Explainability, Uncertainty, and Intervenability AAAI 2024
Distribution-Free Predictive Uncertainty Quantification: Strengths and Limits of Conformal Prediction ICML 2024
Graph Neural Networks: Foundation, Frontiers and Applications KDD 2023
Large-Scale Graph Neural Networks: The Past and New Frontiers KDD 2023
Hyperbolic Graph Neural Networks: A Tutorial on Methods and Applications KDD 2023
Fairness in Graph Machine Learning: Recent Advances and Future Prospectives KDD 2023
Graph Neural Networks: Foundation, Frontiers and Applications AAAI 2023
Temporal Graph Mining for Fraud Detection ICDM 2023
Recent Advances in Bayesian Optimization AAAI 2023
Uncertainty Quantification in Deep Learning KDD 2023
Modeling and Exploiting Data Heterogeneity under Distribution Shifts NeurIPS 2023
Trustworthy Graph Learning: Reliability, Explainability, and Privacy Protection KDD 2022
Algorithmic Fairness on Graphs: Methods and Trends KDD 2022
Automated Learning from Graph-Structured Data AAAI 2022
Fairness in Graph Mining: Metrics, Algorithms, and Applications ICDM 2022
Advances in Bayesian Optimization NeurIPS 2022
Bayesian Optimization: From Foundations to Advanced Topics AAAI 2022
Sampling as First-Order Optimization over a space of probability measures ICML 2022
Graph Representation Learning: Foundations, Methods, Applications and Systems KDD 2021
Toward Explainable Deep Anomaly Detection KDD 2021
The Art of Gaussian Processes: Classical and Contemporary NeurIPS 2021
On Calibration and Out-of-Domain Generalization ICLR 2021
Deep Graph Learning: Foundations, Advances and Applications KDD 2020
Differential Deep Learning on Graphs and its Applications AAAI 2020
Bayesian Deep Learning and a Probabilistic Perspective of Model Construction ICML 2020
Robust Deep Learning Methods for Anomaly Detection. KDD 2020
How to calibrate your neural network classifier: Getting true probabilities from a classification model KDD 2020
Practical Uncertainty Estimation and Out-of-Distribution Robustness in Deep Learning NeurIPS 2020
Graph Representation Learning AAAI 2019
Representation Learning on Graphs and Manifolds ICLR 2019
A Primer on PAC-Bayesian Learning ICML 2019
Deep Learning with Bayesian Principles NeurIPS 2019
Deep Bayesian Mining, Learning and Understanding KDD 2019
Deep Bayesian and Sequential Learning AAAI 2019
Variational Bayes and Beyond: Bayesian Inference for Big Data ICML 2018
Scalable Bayesian Inference NeurIPS 2018
Bayesian Incremental Learning for Deep Neural Networks ICLR 2018
Bayesian Approaches for Blackbox Optimization UAI 2018
Deep Probabilistic Modelling with Gaussian Processes NeurIPS 2017
Non-IID Learning KDD 2017
Bayesian Time Series Modeling: Structured Representations for Scalability ICML 2015
Optimal Algorithms for Learning Bayesian Network Structures UAI 2015
Computational Complexity of Bayesian Networks UAI 2015
Bayesian Posterior Inference in the Big Data Arena: An introduction to probabilistic programming ICML 2014
Approximate Bayesian Computation (ABC) NeurIPS 2013
Linear Programming Relaxations for Graphical Models NeurIPS 2011
Modern Bayesian Nonparametrics NeurIPS 2011